### v1

# Normalized mutual information is a biased measure for classification and community detection

Abstract
Normalized mutual information is widely used as a similarity measure for evaluating the performance of clustering and classification algorithms. In this paper, we show that results returned by the normalized mutual information are biased for two reasons: first, because they ignore the information content of the contingency table and, second, because their symmetric normalization introduces spurious dependence on algorithm output. We introduce a modified version of the mutual information that remedies both of these shortcomings. As a practical demonstration of the importance of using an unbiased measure, we perform extensive numerical tests on a basket of popular algorithms for network community detection and show that one’s conclusions about which algorithm is best are significantly affected by the biases in the traditional mutual information.
I Introduction
A common task in data analysis is the comparison of two different labelings of the same set of objects. How well do demographics predict political affiliation? How accurately do blood tests predict clinical outcomes? How well do clustering algorithms recover known classes of items? The information theoretic measure known as mutual information is commonly applied to answer questions like these, in which an experimental or computational estimate of some kind is compared against a “ground truth.”
Mutual information CT06 is a measure of how easy it is to describe one labeling of a set of objects if we already know another labeling. Specifically, it measures how much less information it takes (in the Shannon sense) to communicate the first labeling if we know the second versus if we do not. As an example, mutual information is commonly used in network science to evaluate the performance of algorithms for network community detection DDDA05 . One takes a network whose community structure is already known and applies a community detection algorithm to it to infer the communities. Then one uses mutual information to compare the output of the algorithm to the known correct communities. Algorithms that consistently achieve high mutual information scores are considered good. We will use this application as an illustrative example later in the paper.
Mutual information has a number of appealing properties as a tool for comparing labelings. It is invariant under permutations of the labels, so that labelings do not have to be aligned before comparison. It also generalizes gracefully to the case where the number of distinct labels is not the same in the two labelings. On the other hand, the mutual information, in its most common form, also has some significant drawbacks, and two in particular that we highlight in this paper. First, it has a bias towards labelings with too many distinct label values. For instance, a community detection algorithm that routinely divides networks into significantly more groups than are present in the ground truth can nonetheless achieve high mutual information scores. Indeed, if the algorithm simply places every node of the network in a group on its own, the mutual information gives this division a maximum score, regardless of the ground truth. In most cases this would be considered an error. A number of approaches for correcting this flaw have been proposed. One can apply direct penalties for incorrect numbers of groups AP17 or subtract correction terms based on the average value of the mutual information over some ensemble of candidate labelings VEB10 ; GA17 or the statistics of the contingency table NCY20 . For reasons we discuss shortly, we favor the latter approach, which leads to the measure known as the reduced mutual information.
The second drawback of the mutual information arises when the measure is normalized, as it commonly is to improve interpretability. The most popular normalization scheme creates a measure that runs between zero and one by dividing the mutual information by the arithmetic mean of the entropies of the two labelings being compared FJ03 , although one can also normalize by the minimum, maximum, or geometric mean of the entropies. As we demonstrate in this paper, however, these normalizations introduce biases into the results by comparison with the unnormalized measure, because the normalization factor depends on the candidate labeling as well as the ground truth. This effect can be large enough to change the ranking of candidate labelings, and we provide examples of this phenomenon.
In order to avoid this latter bias, while still retaining the interpretability of a normalized mutual information score, we propose normalizing the mutual information by the entropy of the ground-truth labeling alone. This removes the source of bias, at the cost of introducing an asymmetry in the normalization. However, given the asymmetric nature of the classification problems we consider, where one or many candidates are compared against a single ground truth, we argue that this is justified. Furthermore, by contrast with the multitude of possible symmetric normalizations, the measure we propose is the unique unbiased way to normalize the mutual information such that a perfect match receives a score of 1.
Both drawbacks of the standard mutual information can be addressed simultaneously by defining an asymmetrically normalized version of the reduced mutual information, and we advocate for this measure in comparing candidate labelings against a known ground truth. In support of this approach we present an extensive comparison of the performance of this and other variants of the mutual information in network community detection tasks, generating a large number of random test networks with known community structure and a variety of structural parameters then attempting to recover the communities using popular community detection algorithms. Within this framework, we find that conclusions about which algorithms perform best are significantly impacted by the choice of mutual information measure, and specifically that algorithms that find too many communities are erroneously favored by traditional measures.
II Mutual information
Mutual information can be thought of in terms of the amount of information it takes to transmit a labeling from one person to another. In general, if we have a set of $N$ possible values for a quantity, such as possible labelings of a set of objects, then one can communicate a single one of them to a receiver by assigning each value a unique binary string and then transmitting the appropriate string. The minimum number of bits needed to do this is
$H=\lceil\log_{2}N\rceil\simeq\log_{2}N\propto\ln N.$ (1)
Conventionally one uses base-2 logarithms in this context, which gives $H$ units of bits, but we use natural logarithms in this paper, since they are more convenient for our purposes. The only difference is an overall factor of $\ln 2$, which will have no effect on our conclusions. Thus, henceforth, all logarithms should be read as natural logs.
Suppose now that we have a labeling or division of $n$ objects into $q$ groups, which we represent as a vector of $n$ integer elements, each with value in the range $1\ldots q$. We assume there to be a ground-truth labeling, which we denote $g$, and a candidate labeling for comparison $c$, generated for instance by some sort of algorithm. The mutual information $I(c;g)$ between the two is the amount of information that can be saved when transmitting the truth $g$ if the receiver already knows the candidate $c$. We can write this information as the total entropy of $g$ minus the conditional entropy:
$I(c;g)=H(g)-H(g|c).$ (2)
It is simple to show that, despite appearances, this definition is symmetric under the interchange of the labelings $c$ and $g$.
We now write down expressions for both entropies in (2), being careful to retain some terms that are often neglected. First consider $H(g)$, which represents the information required to transmit the ground truth on its own. The transmission process has three steps. First, we transmit the number of groups $q_{g}$ in the labeling. The maximum possible value of $q_{g}$ is $n$, so transmitting any particular value requires information $H(q_{g})=\log n$ (see Eq. (1)). Next we transmit a vector $n^{(g)}$ with $q_{g}$ elements $n^{(g)}_{s}$ equal to the number of objects in each group. By definition, the elements of $n^{(g)}$ sum to $n$, and the number of $q_{g}$-element vectors of positive integers that sum to $n$ is $\binom{n-1}{q_{g}-1}$, so the amount of information required to transmit the vector is
$H(n^{(g)}|q_{g})=\log\binom{n-1}{q_{g}-1}.$ (3)
Finally, given $n^{(g)}$, there are $n!/\prod_{s}n_{s}^{(g)}!$ labelings $g$ that have the correct group sizes, so the amount of information needed to identify a particular labeling uniquely is
$H(g|n^{(g)})=\log\frac{n!}{\prod_{s}n_{s}^{(g)}!}.$ (4)
The total cost to transmit $g$ is equal to the sum of the individual costs of the three steps, thus:
$H(g)$ $=H(q_{g})+H(n^{(g)}|q_{g})+H(g|n^{(g)})$
$=\log n+\log\binom{n-1}{q_{g}-1}+\log\frac{n!}{\prod_{s}n_{s}^{(g)}!}.$ (5)
This three-part encoding scheme is not the only one that could be applied to this problem, but it is an efficient one in the common case of a relatively small number of groups $q_{g}\ll n$ with potentially unequal sizes, and it is the one upon which the conventional definition of mutual information is based, which is our focus here.
To calculate the second term in Eq. (2) we use a similar multipart scheme, but one that now takes advantage of the receiver’s knowledge of $c$ when communicating $g$. We first communicate $q_{g}$ and $n^{(g)}$ as before, at the same information cost of $H(q_{g})+H(n^{(g)}|q_{g})=\log n+\log\binom{n-1}{q_{g}-1}$. Then we communicate a contingency table $n^{(cg)}$, a matrix with elements $n_{rs}^{(cg)}$ equal to the number of objects that simultaneously belong to group $r$ in the candidate labeling $c$ and group $s$ in the ground truth $g$. Figure 1 shows an example of a contingency table for two labelings of the nodes in a small network.
Figure 1: An example of a contingency table for two colorings of the same network, one with three colors and one with two. The entries in the $3\times 2$ contingency table $n^{(cg)}$ count the number of nodes that have each combination of colors. The row and column sums $n^{(c)}$ and $n^{(g)}$ of the contingency table count the number of nodes with each color in the two colorings. Note that, although we illustrate the contingency table with an application to a network, the table itself is independent of the network structure.
The contingency table consists of non-negative integer elements and its row and column sums give the sizes of the groups in $c$ and $g$ respectively:
$\sum_{s=1}^{q_{g}}n_{rs}^{(cg)}=n_{r}^{(c)},\qquad\sum_{r=1}^{q_{c}}n_{rs}^{(cg)}=n_{s}^{(g)}.$ (6)
Since the receiver already knows the values of $n^{(c)}$ and $n^{(g)}$ (the former because they know $c$ and the latter because we just transmitted it), only contingency tables with these row and column sums need be considered. The information cost to transmit the contingency table is thus equal to $\log\Omega(n^{(c)},n^{(g)})$, where $\Omega(n^{(c)},n^{(g)})$ is the number of possible tables with the required row and column sums. There is no known general expression for this number, but approximations exist that are good enough for practical purposes DE85 ; BH10 ; JKN22 .
Finally, having transmitted the contingency table, it remains only to transmit the ground-truth labeling itself, where we need consider only those labelings consistent with the contingency table, given the known candidate labeling $c$. The number of such labelings is $\prod_{r}n_{r}^{(c)}!/\prod_{rs}n_{rs}^{(cg)}!$, so the information needed to uniquely identify one of them is
$H(g|n^{(cg)},c)=\log\frac{\prod_{r}n_{r}^{(c)}!}{\prod_{rs}n_{rs}^{(cg)}!}.$ (7)
Putting everything together, the total conditional information is then
$H(g|c)$ $=H(q_{g})+H(n^{(g)}|q_{g})$
$\qquad{}+H(n^{(cg)}|n^{(c)},n^{(g)})+H(g|n^{(cg)},c)$
$=\log n+\log\binom{n-1}{q_{g}-1}$
$\qquad{}+\log\Omega(n^{(c)},n^{(g)})+\log\frac{\prod_{r}n_{r}^{(c)}!}{\prod_{rs}n_{rs}^{(cg)}!}.$ (8)
In typical applications the number of labelings compatible with the contingency table is much smaller than the total number of labelings, and hence this encoding scheme substantially reduces the amount of information needed to transmit the ground truth. This is not the only encoding possible, but it is an efficient one in practice and it is the one upon which the conventional mutual information is based.
Substituting our expressions (5) and (8) for $H(g)$ and $H(g|c)$ into Eq. (2), we now get the mutual information
$I(c;g)$ $=H(g)-H(g|c)$
$=\log\frac{n!\prod_{rs}n_{rs}^{(cg)}!}{\prod_{r}n_{r}^{(c)}!\prod_{s}n_{s}^{(g)}!}-\log\Omega(n^{(c)},n^{(g)}).$ (9)
Note that the terms $\log n$ and $\log\binom{n-1}{q_{g}-1}$ have cancelled and so play no further role. In most treatments the term $\log\Omega(n^{(c)},n^{(g)})$ for the information content of the contingency table is also dropped, since it is subdominant and often small. When this term is retained, the quantity in (9) is known as the reduced mutual information, and there are good reasons to use this full expression in certain circumstances, a point that we investigate in Section II.1. If we follow convention, however, and neglect the contingency table term, then we arrive at the traditional mutual information, which we denote $I_{0}$:
$I_{0}(c;g)=\log\frac{n!\prod_{rs}n_{rs}^{(cg)}!}{\prod_{r}n_{r}^{(c)}!\prod_{s}n_{s}^{(g)}!}.$ (10)
For later convenience, we also define entropies $H_{0}$ for the individual labelings ignoring subdominant terms, which can be thought of as the mutual information between the labelings and themselves:
$H_{0}(c)$ $=I_{0}(c;c)=\log\frac{n!}{\prod_{r}n_{r}^{(c)}!},$ (11)
$H_{0}(g)$ $=I_{0}(g;g)=\log\frac{n!}{\prod_{s}n_{s}^{(g)}!}.$ (12)
Note that these differ from the entropy of Eq. (5), which includes the subdominant terms. Trivially (11) and (12) provide upper bounds on the mutual information, while zero is a lower bound, so
$0\leq I_{0}(c;g)\leq H_{0}(c),H_{0}(g).$ (13)
These bounds will be useful shortly.
A common further step is to approximate the mutual information of Eq. (10) using Stirling’s approximation in the form
$\log n!=n\log n-n+\mathrm{O}(\log(n)),$ (14)
which gives
$I_{0}(c;g)=\sum_{rs}p_{rs}^{(cg)}\log\frac{p_{rs}^{(cg)}}{p_{r}^{(c)}p_{s}^{(g)}}+\mathrm{O}\left(\frac{\log n}{n}\right),$ (15)
where
$p_{rs}^{(cg)}={n_{rs}^{(cg)}\over n},\qquad p_{r}^{(c)}={n_{r}^{(c)}\over n},\qquad p_{s}^{(g)}={n_{s}^{(g)}\over n},$ (16)
and we adopt the convention that $00\log 0=$ 0. Equation (15) is often used as the definition of the mutual information, although technically it is only an approximation, albeit one that is asymptotically correct in the limit of large $n$. (It can also be derived by considering the information cost to transmit the true label of a single object, selected at random, with and without knowledge of the corresponding candidate label.)
II.1 Variants of the mutual information
Equation (10) defines the standard mutual information. In this paper we explore variant definitions that can perform better at the classification tasks we consider. A number of variants have been proposed, of which the simplest in the context of the current presentation is the reduced mutual information $I(c;g)$ of Eq. (9), which is simply the value of the mutual information when one does not neglect the term in $\log\Omega(n^{(c)},n^{(g)})$ that represents the cost of transmitting the contingency table:
$I(c;g)=I_{0}(c;g)-\log\Omega(n^{(c)},n^{(g)}).$ (17)
The moniker “reduced” derives from the fact that the contingency table term is always negative and so reduces the value of the mutual information relative to the conventional definition $I_{0}(c;g)$, but we emphasize that functionally we are simply retaining a term that is usually dropped.
As mentioned previously, there is no general closed-form expression for the number $\Omega(n^{(c)},n^{(g)})$ of contingency tables with given row and column sums, and its numerical computation is #P-hard in general DKM97 and hence intractable for all but the smallest of examples. In practice, therefore, the number must be approximated. In this paper we make use of the “effective columns” approximation of JKN22 , which has good performance over a wide range of situations and a simple closed-form expression:
$\Omega(n^{(c)},n^{(g)})\simeq$
$\binom{n+q_{c}\alpha-1}{q_{c}\alpha-1}^{-1}\prod_{r=1}^{q_{c}}\!\binom{n_{r}^{(c)}+\alpha-1}{\alpha-1}\prod_{s=1}^{q_{g}}\!\binom{n_{s}^{(g)}+q_{c}-1}{q_{c}-1},$ (18)
where
$\alpha=\frac{n^{2}-n+\bigl{[}n^{2}-\sum_{s}(n_{s}^{(g)})^{2}\bigr{]}/q_{c}}{\sum_{s}(n_{s}^{(g)})^{2}-n}.$ (19)
This estimate differs from the one originally used for the reduced mutual information in Newman06b , but we favor it here since it performs better in certain regimes.
To understand the importance of the contingency table term in the mutual information, let us explore how things can go wrong if we neglect it. To take an extreme example, suppose an algorithm simply places every object in a group of its own, which corresponds to the candidate labeling $c=(1,\ldots,n)$. No matter what the ground truth labeling is, this choice of $c$ clearly contains no information about it whatsoever, so we might expect the mutual information to be zero. The contingency table in this case is
$n_{rs}^{(cg)}=\biggl{\{}\begin{array}[]{ll}1&\quad\mbox{if $g_{r}=s$,}\\ 0&\quad\mbox{otherwise,}\\ \end{array}$ (20)
so the conventional mutual information of Eq. (10) is
$I_{0}(c;g)=\log\frac{n!}{\prod_{s}n_{s}^{(g)}!}=H_{0}(g).$ (21)
This answer is as wrong as it could be: we expect the mutual information to take the minimum value of zero, but instead it is equal to the entropy $H_{0}(g)$, which is the maximum possible value since $H_{0}(g)$ is an upper bound as we have said. In other words, $I_{0}(c;g)$ would have us believe that this algorithm which puts every object in its own group tells us everything there is to know about the true labeling $g$, when in fact it tells us nothing at all.
The reason for this result is that in this case the contingency table itself uniquely defines $g$, so neglecting it puts the mutual information in error by an amount equal to the complete information cost of the ground truth. If we include the cost of transmitting the contingency table on the other hand, this erroneous behavior disappears. In this case, we can calculate the number of contingency tables exactly:
$\Omega(n^{(c)},n^{(g)})=\frac{n!}{\prod_{s}n_{s}^{(g)}!},$ (22)
so the reduced mutual information is
$I(c;g)=I_{0}(c;g)-\log\frac{n!}{\prod_{s}n_{s}^{(g)}!}=0,$ (23)
which is now the expected answer.
A related shortcoming of the standard mutual information is that for finite $n$ a random labeling $c$ will have positive mutual information with respect to any ground truth in expectation: because mutual information is non-negative, fluctuations due to randomness in finite systems will produce non-negative values only and hence their average will in general be positive VEB10 ; Zhang15 . This seems counterintuitive: we would expect the average value for a random labeling to be zero.
We can rectify this issue by using another variant of the mutual information, which subtracts off the expected value of the measure, thereby making the average zero by definition. To do this we must first decide how a random labeling is to be defined—over what distribution of candidate labelings are we averaging? A popular choice is to take the uniform distribution over labelings that share the same group sizes $n^{(c)}$ as the actual candidate $c$. This yields the adjusted mutual information of Vinh et al. VEB10 :
$I_{A}(c;g)=I_{0}(c;g)-\langle I_{0}(c;g)\rangle_{\{c|n^{(c)}\}},$ (24)
where the expectation is over the ensemble described. This gives $\langle I_{A}(c;g)\rangle_{\{c|n^{(c)}\}}=0$, so a positive value of the adjusted mutual information indicates a candidate labeling that performs above average within this ensemble.
However, the adjusted mutual information defined in this way, and particularly the choice of ensemble, has some problems NCY20 . The set of labelings with given $n^{(c)}$ heavily favors contingency tables with relatively uniform entries, simply because there are many more labelings that correspond to uniform tables than to non-uniform ones. Real contingency tables, on the other hand, tend to be highly non-uniform, since applications of the mutual information tend to focus on labelings that are at least somewhat similar to the ground truth (which produces a non-uniform table). This means that the average used in the adjusted mutual information puts most of its weight on configurations that are very different from those that occur in reality, making it a poor representation of the actual average mutual information.
A solution to this problem is to perform the average not over all labelings consistent with $n^{(c)}$ but over all contingency tables, which makes non-uniform contingency tables just as likely as uniform ones. As shown in NCY20 , however, adjusting the mutual information in this way simply gives us the reduced mutual information of Eq. (17) again. For this reason we favor the reduced mutual information over the adjusted mutual information: in addition to allowing for the information cost of the contingency table, it corrects sensibly for the nonzero expected value. It is also derived naturally from a fully information-theoretic argument, by contrast with the more ad hoc derivation of the adjusted mutual information. Notwithstanding, in Section III we give results using both reduced and adjusted mutual information, and find fairly similar outcomes in the two cases.
II.2 Normalization of the mutual information
A fundamental difficulty with mutual information as a measure of similarity is that its range of values depends on the particular application, which makes it difficult to say when a value is large or small. Is a mutual information of 10 a large value? Sometimes it is and sometimes it isn’t, depending on the context. To get around this problem one commonly normalizes the mutual information so that it takes a maximum value of 1 when the candidate labeling agrees exactly with the ground truth. There are a number of ways this can be achieved but, as we show here, they are not all equal. In particular, some, including the most popularly used normalization, can result in biased results and should, in our opinion, be avoided. In its place, we propose a novel, unbiased normalized measure.
The most popular normalized measure, commonly referred to simply as the “normalized mutual information,” uses the plain mutual information $I_{0}(c;g)$ as a base measure and normalizes it thus:
$I_{0}^{(S)}(c;g)$ $=\frac{I_{0}(c;g)}{\frac{1}{2}[H_{0}(c)+H_{0}(g)]}$
$=\frac{I_{0}(c;g)}{\frac{1}{2}[I_{0}(c;c)+I_{0}(g;g)]}.$ (25)
This measure has a number of desirable features. Because of the inequalities in (13) its value falls between zero and one. And since both the base measure and the normalization are symmetric under interchange of $c$ and $g$, the normalized measure also retains this symmetry (hence the superscript “$(S)$,” for symmetric).
Figure 2: An example of how normalization can impact which labeling is preferred by a mutual information measure. For the ground-truth labeling $g$ (top), the standard unnormalized mutual information $I_{0}(c,g)$ of Eq. (15) gives different scores to the two candidate labelings $c_{A}$ and $c_{B}$, the former receiving a higher score than the latter. By definition, the asymmetrically normalized mutual information will always agree with the unnormalized measure, but the symmetrically normalized measure may not, as is the case here: in this example the symmetric measure scores candidate $c_{B}$ higher. Note that the network itself plays no role here—it is included only as a visual aid.
Equation (25) is not the only normalization that achieves these goals however. Equation (13) implies that
$I_{0}(c;g)$ $\leq\min(I_{0}(c;c),I_{0}(g;g))$ (26)
$\leq\sqrt{I_{0}(c;c)I_{0}(g;g)}$ (27)
$\leq\max(I_{0}(c;c),I_{0}(g;g)),$ (28)
which gives us three more options for the denominator of the normalized measure. The arithmetic mean in Eq. (25), however, sees the most use by far LF09 ; OLC11 ; YAT16 ; fortunato2016community .
We can extend the notion of symmetric normalization to any other base measure of mutual information $I_{X}(c;g)$ by writing
$I_{X}^{(S)}(c;g)=\frac{I_{X}(c;g)}{\frac{1}{2}[I_{X}(c;c)+I_{X}(g;g)]}.$ (29)
All such measures, however, including the standard measure of Eq. (25), share a crucial shortcoming, namely that the normalization depends on the candidate labeling $c$ and hence that the normalized measure can prefer a different candidate labeling to the base measure purely because of the normalization. Figure 2 shows an example of how this can occur. Two candidate labelings $c_{A}$ and $c_{B}$ are considered for the same ground truth $g$. Under the unnormalized mutual information of Eq. (10), candidate A receives a higher score than candidate B, but under the normalized measure of Eq. (25) the reverse is true. This behavior is due to the difference in entropy between the two candidate divisions, the one on the right having larger entropy than the one on the left, which increases its normalization factor and correspondingly decreases the normalized mutual information.
We contend that in principle the unnormalized measure is more correct on this question, having a direct justification in terms of information theory. The purpose of the normalization is purely to map the values of the measure onto a convenient numerical interval, and should not change outcomes as it does here. Moreover, different symmetric normalizations can produce different results. For instance, if one normalizes by $\max(I_{0}(c;c),I_{0}(g;g))$ in Fig. 2 then candidate $c_{A}$ is favored in all cases.
These issues are unavoidable when using a symmetric normalization scheme. In any such scheme the normalization must depend on both $c$ and $g$ and hence can vary with the candidate labeling. However, if we drop the requirement of symmetry then we can normalize in a straightforward way that avoids these issues. We define the asymmetric normalization of any base measure $I_{X}$ as
$I_{X}^{(A)}(c;g)=\frac{I_{X}(c;g)}{I_{X}(g;g)}.$ (30)
This definition still gives $I_{X}^{(A)}(g;g)=1$, but now the normalization factor in the denominator has no effect on choices between candidate labelings, since it is independent of $c$. In fact, Eq. (30) is the only way to normalize such that $I_{X}^{(A)}(g;g)=1$ while simultaneously ensuring that the preferred candidate is always the same as for the base measure. Thus this measure also removes any ambiguity about how one should perform the normalization.
The downside of this measure is the loss of the symmetry, which is undesirable in some applications. For instance, if one is comparing two candidate labelings directly to one another, rather than to a distinct ground truth, then a symmetric measure may be preferable. Embedding and visualization methods that employ mutual information as a similarity measure also normally demand symmetry PLC17 . In the most common applications of the normalized mutual information, however, where we are evaluating against a ground truth, which is an inherently asymmetric situation, the asymmetric measure makes perfect sense.
Combining the benefits of asymmetric normalization and the reduced mutual information, we advocate in favor of the asymmetrically-normalized reduced mutual information defined by
$I^{(A)}(c;g)=\frac{I(c;g)}{I(g;g)},$ (31)
where
$I(c;g)=I_{0}(c;g)-\log\Omega(n^{(c)},n^{(g)}),$ (32)
with $\Omega(n^{(c)},n^{(g)})$ approximated as in Eq. (18). This measure correctly accounts for the information contained in the contingency table, returns value 0 in expectation when $c$ is uncorrelated with the ground truth and value 1 when $c=g$, and always favors the same labeling as the unnormalized measure.
One potential issue is that at present we are unable to prove a bound $I(c;g)\leq I(g;g)$ analogous to (13), which would ensure that $I^{(A)}(c;g)\leq 1$ always. There is strong numerical and other evidence that such a bound does apply—see Appendix B—but proving it is an open problem.
III Example application:
Community detection
As an example of the performance of our favored measure of mutual information and competing measures, we describe in this section the results of an extensive series of tests in which these measures are used to score the output of various algorithms for network community detection. In these tests we use the popular Lancichinetti-Fortunato-Radicchi (LFR) graph model LFR08 to generate random test networks with known community structure and realistic distributions of node degrees and group sizes, and then attempt to recover that structure using a variety of standard community detection algorithms, quantifying the accuracy of the recovery with six different measures: symmetrically normalized versions of the traditional mutual information, the adjusted mutual information, and the reduced mutual information, and asymmetrically normalized versions of the same. A number of studies have been performed in the past to test the efficacy of community detection algorithms on LFR benchmark networks LF09 ; OLC11 ; YAT16 ; fortunato2016community , but using only the symmetric non-reduced mutual information as a similarity measure. Our results indicate that this measure can produce biased outcomes and we recommend the asymmetric reduced mutual information instead.
The LFR model contains a number of free parameters that control the size of the networks generated, their degree distribution, the distribution of community sizes, and the relative probability of within- and between-group edges. (We give further details on the LFR generative process in Appendix A.) We find that the distributions of degrees and community sizes do not significantly impact the relative performance of the various algorithms tested and that performance differences are driven primarily by the size $n$ of the networks and the mixing parameter $\mu$ that controls the ratio of within/between-group connections, so our tests focus on performance as a function of these variables.
We perform community detection using six well-known algorithms as follows. (We use the implementations found in the igraph library CN06 , except for the inference method, for which we use the graph-tool library P14 .)
  1. 1.
    InfoMap: InfoMap is an information theoretic approach that defines a compression algorithm for encoding a random walk on a network based on which communities the walk passes through RB08 . Different community labelings give rise to more or less efficient compression, as quantified by the so-called map equation, and the labeling with the highest efficiency is considered the best community division.
  2. 2.
    Modularity maximization: Modularity is a quality function for network labelings equal to the fraction of edges within communities minus the expected such fraction if edge positions are randomized while preserving the node degrees. Modularity maximization aims to optimize the modularity over possible labelings and return the labeling with the highest modularity found. Modularity can be maximized using a number of heuristic methods, of which the most popular are agglomerative methods such as the Louvain and Leiden algorithms BGLL08 ; TWV19 , spectral methods Newman06b , and simulated annealing GSA04 ; MAD05 ; RB06a . In our tests we use simulated annealing where computationally feasible and the Leiden algorithm otherwise, these approaches giving the most consistent maximization of the modularity.
  3. 3.
    Modularity with enhanced resolution: Standard modularity maximization is known to suffer from a “resolution limit”—it cannot detect communities smaller than a certain threshold size FB07 . This can be remedied using a variant of modularity that includes a resolution parameter $\gamma$ such that higher values of $\gamma$ cause the algorithm to prefer smaller communities RB06a . Standard modularity maximization corresponds to $\gamma=1$, but for comparison we also conduct tests with $\gamma=10$ using the Leiden algorithm.
  4. 4.
    Statistical inference: Another popular approach to community detection makes use of model fitting and statistical inference. In this context the most commonly fitted model is the degree-corrected stochastic block model KN11a , which can be fitted using Bayesian methods to find the best community division Peixoto17 .
  5. 5.
    Walktrap: Walktrap is an agglomerative algorithm in which initially separate nodes are iteratively combined into progressively larger communities in order from strongest to weakest connections, where strength is quantified in terms of the time for a random walk to reach one node from another PL05 .
  6. 6.
    Labelprop: The label propagation or “labelprop” algorithm initially places every node in its own community then iteratively updates the labels of randomly chosen nodes by majority vote among their network neighbors, breaking ties at random RRS07 .
As described in Section III.2, all of these algorithms perform reasonably well, but, as we will see, the best performers in the context of the current tests are InfoMap and the two variants of modularity maximization.
Figure 3: The colors in each panel in the center of this figure indicate which of three community detection algorithms (InfoMap (red) and modularity maximization with $\gamma=1$ (green) and $\gamma=10$ (blue)—see key on right) are best able to find the known communities in a large set of LFR benchmark networks, according to the six mutual information measures we consider. Mixtures of red, green, and blue denote the proportions of test cases in which each algorithm performs best. Regions in gray indicate parameter values for which no algorithm achieved a positive mutual information score.
III.1 Comparison between variants of the
mutual information
Figure 3 summarizes the relative performance of the various mutual information measures in our tests. In this set of tests we limited ourselves, for the sake of clarity, to the top three community detection algorithms—InfoMap and the two variants of modularity maximization—and measured which of the three returned the best results according to each of our six mutual information measures, as a function of network size $n$ and the mixing parameter $\mu$. Each point in each of the six panels in the figure is color-coded with some mix of red, green, and blue to indicate in what fraction of cases each of the algorithms performs best according to the relevant measure and, as we can see, the results vary significantly among the measures. An experimenter trying to choose the best algorithm would come to substantially different conclusions depending on which measure they use.
One consistent feature of all six mutual information measures is the large red area in each panel of Fig. 3, which represents the region in which InfoMap performs best. Regardless of the measure used, InfoMap is the best performer on networks with low mixing parameter (i.e., strong community structure) and relatively large network size. For higher mixing (weaker structure) or smaller network sizes, modularity maximization does better. Which version of modularity is best, however, depends strongly on the mutual information measure. The traditional symmetrically normalized mutual information (top left panel) mostly favors the version with a high resolution parameter of $\gamma=10$ (blue), but the asymmetric reduced measure for which we advocate (bottom right) favors the version with $\gamma=1$ (green). (The regions colored gray in the figure are those in which no algorithm receives a positive mutual information score and hence all algorithms can be interpreted as failing.)
These results raise significant doubts about the traditional measure. At large $n$ and high mixing $\mu$ the network will have many small communities, and in this regime it is natural that modularity maximization with a high resolution parameter will perform well because it favors large numbers of communities, and all mutual information measures agree on this point. For smaller system sizes or lower mixing, however, the network will have fewer, larger communities and in this regime we find that the $\gamma=10$ algorithm dramatically overestimates the number of communities in the network. With a non-reduced mutual information there is no penalty for doing this, as discussed in Section II.1, which is why the traditional normalized mutual information ranks the $\gamma=10$ algorithm highly in this regime, arguably erroneously. The reduced mutual information, by contrast, prefers the $\gamma=1$ algorithm. The number of groups inferred by each algorithm for each value of the mixing parameter $\mu$ is shown for $n=3200$ in Figure 4, which highlights these findings.
Figure 4: The number of groups inferred by each of the six algorithms in Fig. 3 for LFR benchmark networks with $n=3200$ nodes and a range of values of the mixing parameter $\mu$. The true number of groups is shown in black. The InfoMap algorithm (red) generates an accurate number of groups for values of $\mu$ up to about 0.5, but beyond this point it erroneously places all nodes in a single group. Standard modularity maximization with resolution parameter $\gamma=1$ (green) underestimates the number of groups, presumably because of the resolution limit on the detection of small groups, but not as severely as the number is overestimated when $\gamma=10$ (blue).
An extreme example of the disagreement between measures can be seen in the lower right corner of each plot in Fig. 3, where most measures return a negative value, indicating that community detection has failed altogether. This is expected: for very weak community structure all detection algorithms are expected to show a “detectability threshold” beyond which they are unable to identify any communities DKMZ11a ; Massoulie14 ; MNS15 . The standard normalized mutual information, however, claims to find community structure in this regime using the $\gamma=10$ version of modularity maximization. This occurs because the $\gamma=10$ algorithm finds many small communities and, as discussed in Section II.1, a labeling with many communities, even completely random ones, is accorded a high score by a non-reduced mutual information. This offers a clear reason to avoid the standard measure.
Figure 5: Performance of each of the six community detection algorithms considered here in identifying the known communities in a large set of LFR benchmark networks, as quantified by the asymmetrically normalized reduced information measure $I^{(A)}(c;g)$ proposed in this paper.
The bottom left panel in Fig. 3 shows results for the asymmetrically normalized version of the traditional mutual information, which, perhaps surprisingly, gives even worse results than the symmetric version, with hardly any region in which the $\gamma=1$ version of modularity maximization outperforms the $\gamma=10$ version. The reason for this is interesting: the symmetric normalization of the traditional measure actually acts as an ad hoc correction against the errors introduced by neglecting the information content of the contingency table and hence reduces the amount by which the symmetric measure incorrectly rewards labelings with too many communities. The asymmetric normalization eliminates this correction and hence performs more poorly. The correct solution to this problem, however, is not to use a symmetric normalization, which can bias outcomes in other ways, but rather to adopt a reduced mutual information measure.
Finally, comparing the middle and right-hand columns of Fig. 3, we see that the results for the adjusted and reduced mutual information measures are quite similar in these tests, although there are some differences. In particular, the adjusted measure appears to find more significant structure for higher mixing than the reduced measure. This occurs because, as discussed in Section II.1 and Ref. NCY20 , the adjusted measure encodes the contingency table in a way that is optimized for more uniform tables than the reduced measure, and thus penalizes uniform tables less severely, leading to overestimates of the mutual information in the regime where detection fails. This provides further evidence in favor of using a reduced mutual information measure.
III.2 Comparison between community detection algorithms
Settling on the asymmetrically normalized reduced mutual information as our preferred measure of similarity, we now ask which community detection algorithm or algorithms perform best according to this measure? We have already given away the answer—InfoMap and modularity maximization get the nod—but here we give evidence for that conclusion.
Figure 5 shows results for all six algorithms listed in Section III. Examining the figure we see that in general the best-performing methods are InfoMap, traditional modularity maximization with $\gamma=1$, and the inference method using the degree-corrected stochastic block model. Among the algorithms considered, InfoMap achieves the highest mutual information scores for lower values of the mixing parameter $\mu$ in the LFR model, but fails abruptly as $\mu$ increases, so that beyond a fairly sharp cutoff around $\mu=0.5$ other algorithms do better. InfoMap’s specific failure mode is that it places all nodes in a single community and this behavior can be used as a simple indicator of the failure regime. In this regime one must use another algorithm. Either modularity maximization or inference using the degree-corrected stochastic block model are reasonable options, but modularity has a slight edge, apart from a thin band of intermediate $\mu$ values which, in the interests of simplicity, we choose to ignore. (We discuss some caveats regarding the relationship between the degree-corrected stochastic block model and the LFR benchmark in Appendix A.)
Thus our recommendations for the best community detection algorithm are relatively straightforward, if we base our choice on performance under the LFR benchmark, as is commonly done in this field. If we are in a regime where InfoMap succeeds, meaning it finds more than one community, then one should use InfoMap. If not, one should use standard modularity maximization with $\gamma=1$. That still leaves open the question of how the modularity should be maximized. In our studies we find the best results with simulated annealing, but simulated annealing is computationally expensive. In regimes where it is not feasible, we recommend using the Leiden algorithm instead. (Tests using other computationally efficient maximization schemes, such as the Louvain and spectral algorithms, generally performed less well than the Leiden algorithm.)
IV Conclusions
In this paper we have examined the performance of a range of mutual information measures for comparing labelings of objects in classification, clustering, or community detection applications. We argue that the commonly used normalized mutual information is biased in two ways: (1) because it ignores the information content of the contingency table, which can be large, and (2) because the symmetric normalization it employs introduces spurious dependence on the labeling. We argue in favor of a different measure, an asymmetrically normalized version of the so-called reduced mutual information, which rectifies both of these shortcomings.
To demonstrate the effects of using different mutual information measures, we have presented results of an extensive set of numerical tests on popular network community detection algorithms, as evaluated by the various measures we consider. We find that conclusions about which algorithms are best depend substantially on which measure we use.
Acknowledgements.
This work was supported in part by the US National Science Foundation under grant DMS–2005899 and by computational resources provided by the Advanced Research Computing initiative at the University of Michigan.
Appendix A Benchmarking
In this appendix we give some additional details of our numerical tests.
A.1 Results for the traditional symmetric normalized mutual information
Figure 6: Performance of each of the six community detection algorithms considered here, as quantified by the conventional symmetrically normalized, non-reduced mutual information $I_{0}^{(S)}(c;g)$.
Figure 6 shows the performance of the same six community detection methods as in Fig. 5, but measured using the standard, symmetrically normalized, non-reduced mutual information $I_{0}^{(S)}(c;g)$. By this measure many of the methods appear to perform implausibly well, far beyond the detectability threshold visible in Fig. 5 in the regime where all methods should by rights fail. Note in particular the high scores achieved by the generalized modularity with $\gamma=10$ in this undetectable regime by virtue of the excessive number of groups it generates.
A.2 LFR network generation
The LFR networks we use for benchmarking are generated using the procedure described in LFR08 , which is as follows.
  1. 1.
    Fix the number of nodes $n$ and mixing parameter $\mu$. In our studies we use node counts in the range $n\in[200,51200]$. The parameter $\mu$ controls the relative number of edges within and between communities. For small $\mu$ there are many more edges within communities than between them, which makes the communities relatively easy to detect. But as $\mu$ gets larger there are more edges between communities and detection becomes more difficult in a manner reminiscent of the detectability threshold in the standard stochastic block model DKMZ11a ; DKMZ11b . In our studies we use values of $\mu$ in the range $[0.2,0.8]$.
  2. 2.
    Draw a degree sequence from a power-law distribution with exponent $\tau_{1}$. Many networks have power-law degree distributions, typically with exponents between 2 and 3 Caldarelli07 , and the LFR model exclusively uses power-law distributions. We use $\tau_{1}=2.5$, with average degree $\langle k\rangle=20$ and maximum degree (which depends on graph size) $k_{\text{max}}=n/10$. Empirically, however, our results do not seem to be very sensitive to these choices.
  3. 3.
    Draw a set of community sizes from a power-law distribution with exponent $\tau_{2}$. Many networks are also found to have community sizes that approximately follow a power law, with typical exponents in the range from 1 to 2 Guimera03 ; CNM04 ; PDFV05 ; LFR08 . We use $\tau_{2}=1.5$ with a minimum community size of $s_{\text{min}}=20$ in all cases, while the maximum community size is set to $s_{\text{max}}=\max\left\{n/10,100\right\}$. Again, results were not particularly sensitive to these choices, provided they produce a valid distribution at all.
  4. 4.
    Assign each node to a community randomly, one node at a time, while ensuring that the community chosen is always large enough to support the added node’s intra-community degree, given by $(1-\mu)k$ where $k$ is the total degree.
  5. 5.
    Rewire the edges attached to each node while preserving the node degrees, until the fraction of edges running between nodes in different communities is approximately $\mu$.
The parameter values above are similar to those used for instance in YAT16 . As in that study, we find that algorithm performance is dictated primarily by the parameters $n$ and $\mu$, so it is these parameters that are varied our summary figures.
The LFR model is similar to a special case of the degree-corrected stochastic block model (DCSBM) KN11a , and hence one might expect that inference-based community detection methods employing the latter model would perform optimally on LFR networks. Specifically, in the limit of an infinite number of sampled networks and perfect optimization of each community detection method, the final performance measure for any algorithm is given by the expectation value of the similarity $M(g,h[A])$ between the ground truth LFR partition $g$ and the partition $h[A]$ of the LFR network $A$ inferred using the algorithm, where the expectation is taken over the ensemble $P(A,g|\theta)$ of LFR networks and partitions $A,g$ generated using parameters $\theta$ (meaning $\mu$, $\tau_{1}$, $\tau_{2}$, etc.). By using the LFR benchmark with parameters $\theta$ to compare the performance of community detection algorithms, we are therefore implicitly defining the “best” algorithm to be the one whose corresponding function $h[A]$ optimizes $\sum_{A,g}P(A,g|\theta)M(g,h[A])$. If the similarity measure we choose is the “all or nothing” error function $M(g,c)=\delta(g,c)$, then the optimal community detection algorithm is trivially the one with
$h[A]=\mathop{\textrm{argmax}}\limits_{g}P(A,g|\theta)=\mathop{\textrm{argmax}}\limits_{g}P(g|A,\theta).$ (33)
In other words, the optimal algorithm simply performs maximum a posteriori estimation under the model from which the network was generated. There is no explicit formula for the posterior probability under the LFR model, but to the extent that it is a special case of the DCSBM, we might expect the DCSBM (with appropriate priors) to give optimal results peixoto2021revealing . The LFR model, however, is not precisely a special case of the DCSBM. In particular, the DCSBM normally assumes a uniform distribution over community sizes, where the LRF model assumes a power law. Moreover, we are not using the crude all-or-nothing error function: our entire purpose in this paper is to develop mutual information measures that aggregrate and weigh different modes of error in a sensible fashion. These differences, it appears, are enough to ensure that the DCSBM does not perform the best in our testing.
Regardless, we emphasize that our use of the LFR benchmark in our analysis is simply for consistency with previous studies of network community detection methods LFR08 ; LF09 ; YAT16 . The justification for our proposed similarity measure, on the other hand, is chiefly its theoretical merit over the conventional (symmetric, non-reduced) normalized mutual information, and is independent of the use of the LFR (or any other) benchmark.
Appendix B Upper bound on the
reduced mutual information
In this appendix, we provide evidence for the inequality
$I(c;g)\leq I(g;g),$ (34)
with the exact equality applying if and only if the labelings $c$ and $g$ are identical up to a permutation of the label names. If true, this inequality would imply that the reduced normalized measure $I^{(A)}(c;g)$ was bounded above by 1, as is the unreduced version $I_{0}(c;g)$. We do not at present know of a general proof of (34), but here we present some weaker theoretical results along with strong numerical evidence of its correctness.
First, we note that the only possible violations of (34) would be for small $n$. As $n$ becomes large, the second (reduction) term in Eq. (9) scales as $\log n$, a subleading correction, and hence is negligible for large $n$ compared to the leading mutual information itself which is of order $n$, and the mutual information alone satisfies (34). We have exhaustively checked the following cases with $n$ up to 200 and find (34) to be satisfied in all of them:
$q_{c}=2,\ q_{g}=2$ for $n\leq 200$,
$q_{c}=3,\ q_{g}=2$ for $n\leq 50$,
$q_{c}=4,\ q_{g}=2$ for $n\leq 30$,
$q_{c}=3,\ q_{g}=3$ for $n\leq 30$,
$q_{c}=4,\ q_{g}=3$ for $n\leq 20$.
We also observe zero violations of (9) in any of the (tens of thousands of) randomly generated examples which went into the creation of Fig. 5.
On the theoretical side, (34) implies that $c=g$ is a global maximum of $I(c;g)$, which we are unable to prove, but we can prove the weaker result that it is a local maximum under the mislabeling of any single node, when the reduction term is calculated using Eq. (18). To do this, we first assume, without loss of generality, that a single node with ground truth community label 1 is mislabeled in the candidate partition $c$ as being in community 2. This implies that $n_{1}^{(c)}=n_{1}^{(g)}-1$, $n_{2}^{(c)}=n_{2}^{(g)}+1$, and $n_{r}^{(c)}=n_{r}^{(g)}$ for all $r=3\ldots q_{g}$. The contingency table will then be nearly diagonal, with $(n_{1}^{(g)}-1,n_{2}^{(g)},\ldots)$ on the diagonal and a single nonzero off-diagonal element $n_{12}^{(cg)}=1$. The decrease in the standard mutual information is then given by
$I_{0}(g;g)-I_{0}(g;c)=\log\frac{n!}{\prod_{s}n_{s}^{(g)}!}$
$\quad{}-\log\frac{n!1!(n_{1}^{(g)}-1)!\prod_{s=2}^{q_{g}}n_{s}^{(g)}!}{\left(\prod_{s}n_{s}^{(g)}!\right)\left(\prod_{s=3}^{q_{g}}n_{s}^{(g)}!\right)(n_{1}^{(g)}-1)!(n_{2}^{(g)}+1)!}$
$=\log(n_{2}^{(g)}+1).$ (35)
At the same time, the reduction term $\log\Omega$ also decreases, by an amount that can be estimated from Eq. (18) thus:
$\log$ $\Omega(n^{(g)},n^{(g)})-\log\Omega(n^{(g)},n^{(c)})=\log\prod_{r=1}^{q_{g}}\frac{\binom{n_{r}^{(g)}+\alpha-1}{\alpha-1}}{\binom{n_{r}^{(c)}+\alpha-1}{\alpha-1}}$
$=\log\frac{(n_{1}^{(g)}+\alpha+1)(n_{2}^{(g)}+1)}{n_{1}^{(g)}(n_{2}^{(g)}+\alpha)}.$ (36)
Combining Eqs. (35) and (36), we now find that the change in the RMI incurred from the mislabeling of a single node is
$I(g;g)-I(g;c)=\log\frac{n_{1}^{(g)}(n_{2}^{(g)}+\alpha)}{n_{1}^{(g)}+\alpha-1}>0,$ (37)
where the inequality holds so long as $n_{1}^{(g)},n_{2}^{(g)},\alpha\geq 1$. Thus, under the estimate of $\Omega(n^{(g)},n^{(c)})$ given by Eq. (18), $g=c$ is a local maximum of the reduced mutual information.

### v2

@@
 Normalized mutual information is widely used as a similarity measure for evaluating the performance of clustering and classification algorithms.
-In this paper, we show that results returned by the normalized mutual information are biased for two reasons: first, because they ignore the information content of the contingency table and, second, because their symmetric normalization introduces spurious dependence on algorithm output.
+In this paper, we argue that results returned by the normalized mutual information are biased for two reasons: first, because they ignore the information content of the contingency table and, second, because their symmetric normalization introduces spurious dependence on algorithm output.
 We introduce a modified version of the mutual information that remedies both of these shortcomings.
@@
 How accurately do blood tests predict clinical outcomes?
-How well do clustering algorithms recover known classes of items?
 The information theoretic measure known as mutual information is commonly applied to answer questions like these, in which an experimental or computational estimate of some kind is compared against a “ground truth.”
-Mutual information CT06 is a measure of how easy it is to describe one labeling of a set of objects if we already know another labeling.
-Specifically, it measures how much less information it takes (in the Shannon sense) to communicate the first labeling if we know the second versus if we do not.
-As an example, mutual information is commonly used in network science to evaluate the performance of algorithms for network community detection DDDA05 .
+Mutual information [1] is a measure of how efficiently we can describe one labeling of the set of objects if we already know another labeling.
+Specifically, it measures how much less information it takes to communicate the first labeling if we know the second versus if we do not.
+As an example, mutual information is commonly used in network science to evaluate the performance of algorithms for network community detection [2].
 One takes a network whose community structure is already known and applies a community detection algorithm to it to infer the communities.
@@
 Mutual information has a number of appealing properties as a tool for comparing labelings.
-It is invariant under permutations of the labels, so that labelings do not have to be aligned before comparison.
-It also generalizes gracefully to the case where the number of distinct labels is not the same in the two labelings.
-On the other hand, the mutual information, in its most common form, also has some significant drawbacks, and two in particular that we highlight in this paper.
+It is invariant under permutations of the labels, so that, for example, two divisions of a network into communities that differ only in how the communities are labeled will be correctly identified as being the same.
+It also returns sensible results in cases where the number of distinct label values is not the same in the two labelings.
+However, the standard mutual information measure also has some significant shortcomings, and two in particular that we highlight in this paper.
 First, it has a bias towards labelings with too many distinct label values.
 For instance, a community detection algorithm that routinely divides networks into significantly more groups than are present in the ground truth can nonetheless achieve high mutual information scores.
-Indeed, if the algorithm simply places every node of the network in a group on its own, the mutual information gives this division a maximum score, regardless of the ground truth.
-In most cases this would be considered an error.
 A number of approaches for correcting this flaw have been proposed.
-One can apply direct penalties for incorrect numbers of groups AP17 or subtract correction terms based on the average value of the mutual information over some ensemble of candidate labelings VEB10 ; GA17 or the statistics of the contingency table NCY20 .
+One can apply direct penalties for incorrect numbers of groups [3] or subtract correction terms based on the average value of the mutual information over some ensemble of candidate labelings [4, 5] or on the statistics of the contingency table [6, 7].
 For reasons we discuss shortly, we favor the latter approach, which leads to the measure known as the reduced mutual information.
 The second drawback of the mutual information arises when the measure is normalized, as it commonly is to improve interpretability.
-The most popular normalization scheme creates a measure that runs between zero and one by dividing the mutual information by the arithmetic mean of the entropies of the two labelings being compared FJ03 , although one can also normalize by the minimum, maximum, or geometric mean of the entropies.
+The most popular normalization scheme creates a measure that runs between zero and one by dividing the mutual information by the arithmetic mean of the entropies of the two labelings being compared [8], although one can also normalize by the minimum, maximum, or geometric mean of the entropies.
 As we demonstrate in this paper, however, these normalizations introduce biases into the results by comparison with the unnormalized measure, because the normalization factor depends on the candidate labeling as well as the ground truth.
-This effect can be large enough to change the ranking of candidate labelings, and we provide examples of this phenomenon.
-In order to avoid this latter bias, while still retaining the interpretability of a normalized mutual information score, we propose normalizing the mutual information by the entropy of the ground-truth labeling alone.
-This removes the source of bias, at the cost of introducing an asymmetry in the normalization.
-However, given the asymmetric nature of the classification problems we consider, where one or many candidates are compared against a single ground truth, we argue that this is justified.
-Furthermore, by contrast with the multitude of possible symmetric normalizations, the measure we propose is the unique unbiased way to normalize the mutual information such that a perfect match receives a score of 1.
-Both drawbacks of the standard mutual information can be addressed simultaneously by defining an asymmetrically normalized version of the reduced mutual information, and we advocate for this measure in comparing candidate labelings against a known ground truth.
-In support of this approach we present an extensive comparison of the performance of this and other variants of the mutual information in network community detection tasks, generating a large number of random test networks with known community structure and a variety of structural parameters then attempting to recover the communities using popular community detection algorithms.
-Within this framework, we find that conclusions about which algorithms perform best are significantly impacted by the choice of mutual information measure, and specifically that algorithms that find too many communities are erroneously favored by traditional measures.
+This effect can be large enough to change scientific conclusions, and we provide examples of this phenomenon.
+In order to avoid this latter bias, while still retaining the interpretability of a normalized mutual information measure, in this paper we propose normalizing by the entropy of the ground-truth labeling alone.
+This removes the source of bias but introduces an asymmetry in the normalization.
+At first sight this asymmetry might seem undesirable, and previous authors have gone to some lengths to avoid it.
+Here, however, we argue that it is not only justified but even desirable, for several reasons.
+First, many of the classification problems we consider are unaffected by the asymmetry, since they involve the comparison of one or more candidate labelings against a single, unchanging ground truth.
+Moreover, by contrast with the multitude of possible symmetric normalizations, the asymmetric measure we propose is the unique unbiased way to normalize the mutual information such that a perfect match receives a score of 1.
+But more broadly, one can argue that an asymmetric measure can be more informative than the conventional symmetric one.
+Consider for instance the common situation where one labeling is a fine-grained version of the other.
+We might for example label individuals by the country they live in on the one hand and by the town or city on the other.
+Then the detailed labeling tells us everything there is to know about the coarse-grained one, but the reverse is not true: telling you the city fixes the country, but not vice versa.
+Thus one could argue that the mutual information between the two should be asymmetric even before normalization—labeling A contains more information about labeling B than B does about A—and this type of asymmetry will be a key feature of the measures we study.
+Both drawbacks of the standard mutual information—bias towards too many groups and dependence of the normalization on the candidate labeling—can be addressed simultaneously by using an asymmetric normalized reduced mutual information as defined in this paper.
+In support of this approach we present an extensive comparison of the performance of this and other variants of the mutual information in network community detection tasks, generating a large number of random test networks with known community structure and a variety of structural parameters, and then attempting to recover the communities using popular community detection algorithms.
+Within this framework, we find that conclusions about which algorithms perform best are significantly impacted by the choice of mutual information measure, and specifically that traditional measures erroneously favor algorithms that find too many communities, but our proposed measure does not.
+Code implementing our measure may be found at https://github.com/maxjerdee/reduced_mutual_information.
 II Mutual information
@@
 The only difference is an overall factor of $\ln 2$, which will have no effect on our conclusions.
-Thus, henceforth, all logarithms should be read as natural logs.
-Suppose now that we have a labeling or division of $n$ objects into $q$ groups, which we represent as a vector of $n$ integer elements, each with value in the range $1\ldots q$.
+Henceforth, all logarithms should be read as natural logs.
+Suppose that we have a labeling or division of $n$ objects into $q$ groups, which we represent as a vector of $n$ integer elements, each with value in the range $1\ldots q$.
 We assume there to be a ground-truth labeling, which we denote $g$, and a candidate labeling for comparison $c$, generated for instance by some sort of algorithm.
 The mutual information $I(c;g)$ between the two is the amount of information that can be saved when transmitting the truth $g$ if the receiver already knows the candidate $c$.
-We can write this information as the total entropy of $g$ minus the conditional entropy:
+We can write this information as the total entropy of $g$ minus the conditional entropy of $g$ given $c$:
 $I(c;g)=H(g)-H(g|c).$ (2)
-It is simple to show that, despite appearances, this definition is symmetric under the interchange of the labelings $c$ and $g$.
-We now write down expressions for both entropies in (2), being careful to retain some terms that are often neglected.
+Loosely, we say that $I(c;g)$ measures “how much $c$ tells us about $g$.”
+We now write down expressions for both entropies in (2), but there are some subtleties involved in doing so, because the expressions depend on how the transmitted information is encoded and also because achieving correct results requires us to carefully retain some terms that are often neglected.
+Our presentation follows that of [7].
 First consider $H(g)$, which represents the information required to transmit the ground truth on its own.
 The transmission process has three steps.
-First, we transmit the number of groups $q_{g}$ in the labeling.
-The maximum possible value of $q_{g}$ is $n$, so transmitting any particular value requires information $H(q_{g})=\log n$ (see Eq.
-(1)).
-Next we transmit a vector $n^{(g)}$ with $q_{g}$ elements $n^{(g)}_{s}$ equal to the number of objects in each group.
-By definition, the elements of $n^{(g)}$ sum to $n$, and the number of $q_{g}$-element vectors of positive integers that sum to $n$ is $\binom{n-1}{q_{g}-1}$, so the amount of information required to transmit the vector is
-$H(n^{(g)}|q_{g})=\log\binom{n-1}{q_{g}-1}.$ (3)
-Finally, given $n^{(g)}$, there are $n!/\prod_{s}n_{s}^{(g)}!$ labelings $g$ that have the correct group sizes, so the amount of information needed to identify a particular labeling uniquely is
+We first transmit the number of groups $q_{g}$ in the labeling, followed by a vector $n^{(g)}$ with $q_{g}$ elements $n^{(g)}_{s}$ equal to the number of objects in each group.
+Third, we transmit our final labeling $g$.
+If we apply Eq.
+(1) to calculate the cost of the first step, since the maximum possible value of $q_{g}$ is $n$, transmitting any particular value requires information $H(q_{g})=\log n$.
+Likewise for the third step, given $n^{(g)}$, there are $n!/\prod_{s}n_{s}^{(g)}!$ labelings $g$ that have the correct group sizes, so the amount of information needed to identify a particular labeling uniquely is
 $H(g|n^{(g)})=\log\frac{n!}{\prod_{s}n_{s}^{(g)}!}.$ (3)
-The total cost to transmit $g$ is equal to the sum of the individual costs of the three steps, thus:
-$H(g)$ $=H(q_{g})+H(n^{(g)}|q_{g})+H(g|n^{(g)})$
-$=\log n+\log\binom{n-1}{q_{g}-1}+\log\frac{n!}{\prod_{s}n_{s}^{(g)}!}.$ (5)
-This three-part encoding scheme is not the only one that could be applied to this problem, but it is an efficient one in the common case of a relatively small number of groups $q_{g}\ll n$ with potentially unequal sizes, and it is the one upon which the conventional definition of mutual information is based, which is our focus here.
-To calculate the second term in Eq.
-(2) we use a similar multipart scheme, but one that now takes advantage of the receiver’s knowledge of $c$ when communicating $g$.
-We first communicate $q_{g}$ and $n^{(g)}$ as before, at the same information cost of $H(q_{g})+H(n^{(g)}|q_{g})=\log n+\log\binom{n-1}{q_{g}-1}$.
-Then we communicate a contingency table $n^{(cg)}$, a matrix with elements $n_{rs}^{(cg)}$ equal to the number of objects that simultaneously belong to group $r$ in the candidate labeling $c$ and group $s$ in the ground truth $g$.
+The middle step, transmission of the group sizes, is more delicate: here some optimization is possible that allows us to achieve a better measure of information content.
+Generally when transmitting a sequence of $n$ values with unequal probabilities such that value $r$ occurs $n_{r}$ times, the information cost is given by the generalization of Eq.
+(3) thus:
+$\log{n!\over\prod_{r}n_{r}!}$ $\simeq n\log n-n-\sum_{r}(n_{r}\log n_{r}-n_{r})$
+$=-\sum_{r}n_{r}\log p_{r},$ (4)
+where $p_{r}=n_{r}/n$ is the probability of value $r$ and we have approximated the factorials using Stirling’s formula.
+Equation 4 tells us that the information cost to transmit the value $r$ is simply $-\log p_{r}$.
+If the $N$ possible values are equally likely to appear and $p_{r}=1/N$, we recover the information cost of Eq.
+(1), but if they are not we can take advantage of that fact to create more efficient encodings.
+The transmission of $n^{(g)}$ benefits from such an efficient encoding.
+We do not in general know the distribution of the elements of $n^{(g)}$, but we know they that may be—and often are—heterogeneous.
+As shown in [7], an effective approach is to assume a Dirichlet-multinomial distribution with concentration parameter $\alpha_{g}>0$
+$P(n^{(g)}|q_{g},\alpha_{g})=$ (5)
+$\qquad\binom{n+q_{g}\alpha_{g}-1}{q_{g}\alpha_{g}-1}^{-1}\prod_{r=1}^{q_{g}}\binom{n_{r}^{(g)}+\alpha_{g}-1}{\alpha_{g}-1},$
+where for non-integer $\alpha_{g}$ we generalize the binomial coefficient in the obvious way:
+$\Bigl{(}{n\atop k}\Bigr{)}={\frac{\Gamma(n+1)}{\Gamma(k+1)\Gamma(n-k+1)}},$ (6)
+with $\Gamma(x)$ being the standard gamma function.
+If $\alpha_{g}=1$, the Dirichlet-multinomial distribution is uniform over all $q_{g}$-vectors $n^{(g)}$ of non-negative integers that sum to $n$.
+If $10<\alpha_{g}<$ 1 the distribution favors vectors with more extreme elements, while for $\alpha_{g}>1$ it favors vectors with more homogeneous elements.
+With this choice the resulting information cost to transmit $n^{(g)}$ is
+$H($ $n^{(g)}|q_{g},\alpha_{g})=-\log P(n^{(g)}|q_{g},\alpha_{g})$ (7)
+$=\log\binom{n+q_{g}\alpha_{g}-1}{q_{g}\alpha_{g}-1}-\sum_{r=1}^{q_{g}}\log\binom{n_{r}^{(g)}+\alpha_{g}-1}{\alpha_{g}-1}.$
+The optimal encoding for transmitting $n^{(g)}$ within the Dirichlet-multinomial family is given by the minimum of this expression with respect to $\alpha_{g}$, which is also equivalent to simply maximizing $P(n^{(g)}|\alpha_{g})$, i.e., to finding the maximum-likelihood value of $\alpha_{g}$.
+Note that we omit the cost of transmitting the parameter $\alpha_{g}$ itself since it will ultimately cancel out of the expression for the mutual information as described in [7].
+The total cost to transmit $g$ is now equal to the sum of the individual costs of the three steps, thus:
+$H(g)$ $=H(q_{g})+H(n^{(g)}|q_{g},\alpha_{g})+H(g|n^{(g)})$
+$=\log n+\log\frac{n!}{\prod_{s}n_{s}^{(g)}!}$ (8)
+$+\log\binom{n+q_{g}\alpha_{g}-1}{q_{g}\alpha_{g}-1}-\sum_{r=1}^{q_{g}}\log\binom{n_{r}^{(g)}+\alpha_{g}-1}{\alpha_{g}-1}$
+This three-part encoding scheme is not the only one that can be applied to this problem, but it is an efficient one in the common case of a relatively small number of groups $q_{g}\ll n$ with potentially unequal sizes, and it reduces to the conventional expression for $H(g)$ when subleading terms are neglected.
+It is an important part of our argument here, however, that the subleading terms be retained and carried through to the calculation of the mutual information.
+For the calculation of $H(g|c)$, the second term in Eq.
+(2), we use a similar three-part encoding, but one that takes advantage of the receiver’s knowledge of $c$ when communicating $g$.
+We first communicate $q_{g}$ as before, at the same information cost of $H(q_{g})=\log n$.
+In place of the vector $n^{(g)}$ we then communicate a contingency table $n^{(gc)}$, a matrix with elements $n_{rs}^{(gc)}$ equal to the number of objects that simultaneously belong to group $r$ in the candidate labeling $c$ and group $s$ in the ground truth $g$.
 Figure 1 shows an example of a contingency table for two labelings of the nodes in a small network.
 Figure 1: An example of a contingency table for two colorings of the same network, one with three colors and one with two.
-The entries in the $3\times 2$ contingency table $n^{(cg)}$ count the number of nodes that have each combination of colors.
-The row and column sums $n^{(c)}$ and $n^{(g)}$ of the contingency table count the number of nodes with each color in the two colorings.
+The entries in the $3\times 2$ contingency table $n^{(gc)}$ count the number of nodes that have each combination of colors.
+The row and column sums $n^{(g)}$ and $n^{(c)}$ of the contingency table count the number of nodes with each color in the two colorings.
 Note that, although we illustrate the contingency table with an application to a network, the table itself is independent of the network structure.
-The contingency table consists of non-negative integer elements and its row and column sums give the sizes of the groups in $c$ and $g$ respectively:
-$\sum_{s=1}^{q_{g}}n_{rs}^{(cg)}=n_{r}^{(c)},\qquad\sum_{r=1}^{q_{c}}n_{rs}^{(cg)}=n_{s}^{(g)}.$ (6)
-Since the receiver already knows the values of $n^{(c)}$ and $n^{(g)}$ (the former because they know $c$ and the latter because we just transmitted it), only contingency tables with these row and column sums need be considered.
-The information cost to transmit the contingency table is thus equal to $\log\Omega(n^{(c)},n^{(g)})$, where $\Omega(n^{(c)},n^{(g)})$ is the number of possible tables with the required row and column sums.
-There is no known general expression for this number, but approximations exist that are good enough for practical purposes DE85 ; BH10 ; JKN22 .
+Our process for transmitting the contingency table is analogous to the one we use for transmitting $n^{(g)}$.
+Since the labeling $c$ is known, the column sums $n^{(c)}$ of the contingency table are known as well, and we assume each column to be drawn from a Dirichlet-multinomial distribution with the appropriate sum.
+We use the same concentration parameter $\alpha_{g|c}$ for every column, but the columns are otherwise independent.
+If we denote column $s$ by $n_{\cdot s}^{(gc)}$, then the information cost to transmit the entire contingency table is
+$H($ $n^{(gc)}|n^{(c)},q_{g},\alpha_{g|c})=\sum_{s=1}^{q_{c}}H(n_{\cdot s}^{(gc)}|n_{s}^{(c)},q_{g},\alpha_{g|c})$
+$=\sum_{s=1}^{q_{c}}\biggl{[}\log\binom{n_{s}^{(c)}+q_{g}\alpha_{g|c}-1}{q_{g}\alpha_{g|c}-1}$
+$\hskip 60.00009pt-\sum_{r=1}^{q_{g}}\log\binom{n_{rs}^{(gc)}+\alpha_{g|c}-1}{\alpha_{g|c}-1}\biggr{]}.$ (9)
+This formulation allows for very efficient transmission of the contingency table under commonly occurring conditions.
+Consider for instance the special (but not implausible) case where $g=c$, so that $n^{(gc)}$ is diagonal.
+In this case the optimal concentration parameter is $\alpha_{g|c}=0$, which restricts the possible contingency tables to those for which each column $n_{\cdot s}^{(gc)}$ has exactly one nonzero entry.
+Given that we know the column sums, this leaves only the positions of the nonzero entries within each column to be transmitted, which results in near-optimal performance, a decisive improvement over previous reductions of the mutual information [7].
+Most practical applications will not have $c=g$ precisely, but the two labelings are often similar and similarly impressive improvements in performance can be realized in such cases as well.
 Finally, having transmitted the contingency table, it remains only to transmit the ground-truth labeling itself, where we need consider only those labelings consistent with the contingency table, given the known candidate labeling $c$.
-The number of such labelings is $\prod_{r}n_{r}^{(c)}!/\prod_{rs}n_{rs}^{(cg)}!$, so the information needed to uniquely identify one of them is
-$H(g|n^{(cg)},c)=\log\frac{\prod_{r}n_{r}^{(c)}!}{\prod_{rs}n_{rs}^{(cg)}!}.$ (7)
-Putting everything together, the total conditional information is then
-$H(g|c)$ $=H(q_{g})+H(n^{(g)}|q_{g})$
-$\qquad{}+H(n^{(cg)}|n^{(c)},n^{(g)})+H(g|n^{(cg)},c)$
-$=\log n+\log\binom{n-1}{q_{g}-1}$
-$\qquad{}+\log\Omega(n^{(c)},n^{(g)})+\log\frac{\prod_{r}n_{r}^{(c)}!}{\prod_{rs}n_{rs}^{(cg)}!}.$ (8)
-In typical applications the number of labelings compatible with the contingency table is much smaller than the total number of labelings, and hence this encoding scheme substantially reduces the amount of information needed to transmit the ground truth.
-This is not the only encoding possible, but it is an efficient one in practice and it is the one upon which the conventional mutual information is based.
-Substituting our expressions (5) and (8) for $H(g)$ and $H(g|c)$ into Eq. (2), we now get the mutual information
+The number of such labelings is $\prod_{r}n_{r}^{(c)}!/\prod_{rs}n_{rs}^{(gc)}!$, so the information needed to uniquely identify one of them is
+$H(g|n^{(gc)},c)=\log\frac{\prod_{r}n_{r}^{(c)}!}{\prod_{rs}n_{rs}^{(gc)}!}.$ (10)
+In typical applications the number of labelings compatible with the contingency table is much smaller than the total number of labelings, and hence this encoding substantially reduces the amount of information needed to transmit the ground truth.
+It is not the only encoding possible, but it is an efficient one in practice and it is the one upon which the conventional mutual information is based.
+Collecting all our terms, the total conditional information can now be written
+$H(g|c)$ $=H(q_{g})+H(n^{(gc)}|n^{(c)},q_{g},\alpha_{g|c})+H(g|n^{(cg)},c).$ (11)
+And substituting (8) and (11) into Eq. (2), we get a complete expression for the mutual information
 $I(c;g)$ $=H(g)-H(g|c)$
-$=\log\frac{n!\prod_{rs}n_{rs}^{(cg)}!}{\prod_{r}n_{r}^{(c)}!\prod_{s}n_{s}^{(g)}!}-\log\Omega(n^{(c)},n^{(g)}).$ (9)
-Note that the terms $\log n$ and $\log\binom{n-1}{q_{g}-1}$ have cancelled and so play no further role.
-In most treatments the term $\log\Omega(n^{(c)},n^{(g)})$ for the information content of the contingency table is also dropped, since it is subdominant and often small.
-When this term is retained, the quantity in (9) is known as the reduced mutual information, and there are good reasons to use this full expression in certain circumstances, a point that we investigate in Section II.1.
-If we follow convention, however, and neglect the contingency table term, then we arrive at the traditional mutual information, which we denote $I_{0}$:
-$I_{0}(c;g)=\log\frac{n!\prod_{rs}n_{rs}^{(cg)}!}{\prod_{r}n_{r}^{(c)}!\prod_{s}n_{s}^{(g)}!}.$ (10)
+$=\log\frac{n!\prod_{rs}n_{rs}^{(gc)}!}{\prod_{r}n_{r}^{(c)}!\prod_{s}n_{s}^{(g)}!}+\log\binom{n+q_{g}\alpha_{g}-1}{q_{g}\alpha_{g}-1}$
+$\qquad-\sum_{r=1}^{q_{g}}\log\binom{n_{r}^{(g)}+\alpha_{g}-1}{\alpha_{g}-1}$
+$\qquad-\sum_{s=1}^{q_{c}}\log\binom{n_{s}^{(c)}+q_{g}\alpha_{g|c}-1}{q_{g}\alpha_{g|c}-1}$
+$\qquad+\sum_{r=1}^{q_{g}}\sum_{s=1}^{q_{c}}\log\binom{n_{rs}^{(gc)}+\alpha_{g|c}-1}{\alpha_{g|c}+1}.$ (12)
+In most treatments, only the first of these terms is retained, since the others are subdominant and often small [6, 7].
+When all the terms are retained, however, the quantity (12) is known as the reduced mutual information and there are good reasons to use this full expression in certain circumstances, a point that we investigate in Section II.1.
+If we follow convention, however, and keep only the leading term, then we arrive at the traditional mutual information, which we denote $I_{0}$:
+$I_{0}(c;g)=\log\frac{n!\prod_{rs}n_{rs}^{(gc)}!}{\prod_{r}n_{r}^{(c)}!\prod_{s}n_{s}^{(g)}!}.$ (13)
 For later convenience, we also define entropies $H_{0}$ for the individual labelings ignoring subdominant terms, which can be thought of as the mutual information between the labelings and themselves:
@@
 $H_{0}(g)$ $=I_{0}(g;g)=\log\frac{n!}{\prod_{s}n_{s}^{(g)}!}.$ (15)
-Note that these differ from the entropy of Eq. (5), which includes the subdominant terms. Trivially (11) and (12) provide upper bounds on the mutual information, while zero is a lower bound, so
+Note that these differ from the entropy of Eq. (8), which includes subdominant terms. Trivially (14) and (15) provide upper bounds on the mutual information, while zero is a lower bound, so
 $0\leq I_{0}(c;g)\leq H_{0}(c),H_{0}(g).$ (16)
@@
 which gives
-$I_{0}(c;g)=\sum_{rs}p_{rs}^{(cg)}\log\frac{p_{rs}^{(cg)}}{p_{r}^{(c)}p_{s}^{(g)}}+\mathrm{O}\left(\frac{\log n}{n}\right),$ (15)
+$I_{0}(c;g)=n\sum_{rs}p_{rs}^{(gc)}\log\frac{p_{rs}^{(gc)}}{p_{r}^{(c)}p_{s}^{(g)}}+\mathrm{O}\left(\log n\right),$ (18)
 where
-$p_{rs}^{(cg)}={n_{rs}^{(cg)}\over n},\qquad p_{r}^{(c)}={n_{r}^{(c)}\over n},\qquad p_{s}^{(g)}={n_{s}^{(g)}\over n},$ (16)
+$p_{rs}^{(gc)}={n_{rs}^{(gc)}\over n},\qquad p_{r}^{(c)}={n_{r}^{(c)}\over n},\qquad p_{s}^{(g)}={n_{s}^{(g)}\over n},$ (19)
 and we adopt the convention that $00\log 0=$ 0.
@@
 (It can also be derived by considering the information cost to transmit the true label of a single object, selected at random, with and without knowledge of the corresponding candidate label.)
-II.1 Variants of the mutual information
+II.1 Reduced mutual information
 Equation (13) defines the standard mutual information.
-In this paper we explore variant definitions that can perform better at the classification tasks we consider.
+In this paper we explore variant definitions that can perform better at the clustering tasks we consider.
 A number of variants have been proposed, of which the simplest in the context of the current presentation is the reduced mutual information $I(c;g)$ of Eq.
-(9), which is simply the value of the mutual information when one does not neglect the term in $\log\Omega(n^{(c)},n^{(g)})$ that represents the cost of transmitting the contingency table:
-$I(c;g)=I_{0}(c;g)-\log\Omega(n^{(c)},n^{(g)}).$ (17)
-The moniker “reduced” derives from the fact that the contingency table term is always negative and so reduces the value of the mutual information relative to the conventional definition $I_{0}(c;g)$, but we emphasize that functionally we are simply retaining a term that is usually dropped.
-As mentioned previously, there is no general closed-form expression for the number $\Omega(n^{(c)},n^{(g)})$ of contingency tables with given row and column sums, and its numerical computation is #P-hard in general DKM97 and hence intractable for all but the smallest of examples.
-In practice, therefore, the number must be approximated.
-In this paper we make use of the “effective columns” approximation of JKN22 , which has good performance over a wide range of situations and a simple closed-form expression:
-$\Omega(n^{(c)},n^{(g)})\simeq$
-$\binom{n+q_{c}\alpha-1}{q_{c}\alpha-1}^{-1}\prod_{r=1}^{q_{c}}\!\binom{n_{r}^{(c)}+\alpha-1}{\alpha-1}\prod_{s=1}^{q_{g}}\!\binom{n_{s}^{(g)}+q_{c}-1}{q_{c}-1},$ (18)
-where
-$\alpha=\frac{n^{2}-n+\bigl{[}n^{2}-\sum_{s}(n_{s}^{(g)})^{2}\bigr{]}/q_{c}}{\sum_{s}(n_{s}^{(g)})^{2}-n}.$ (19)
-This estimate differs from the one originally used for the reduced mutual information in Newman06b , but we favor it here since it performs better in certain regimes.
-To understand the importance of the contingency table term in the mutual information, let us explore how things can go wrong if we neglect it.
+(12), which is simply the value of the mutual information when one does not neglect the subleading information cost of transmitting the group sizes and the contingency table.
+To understand the importance of the subleading terms let us explore how things can go wrong if we neglect them.
 To take an extreme example, suppose an algorithm simply places every object in a group of its own, which corresponds to the candidate labeling $c=(1,\ldots,n)$.
-No matter what the ground truth labeling is, this choice of $c$ clearly contains no information about it whatsoever, so we might expect the mutual information to be zero.
+No matter what the ground truth is, this choice of $c$ clearly contains no information about it whatsoever, so we might expect the mutual information to be zero.
+This, however, is not what we find.
 The contingency table in this case is
-$n_{rs}^{(cg)}=\biggl{\{}\begin{array}[]{ll}1&\quad\mbox{if $g_{r}=s$,}\\ 0&\quad\mbox{otherwise,}\\ \end{array}$ (20)
+$n_{rs}^{(gc)}=\biggl{\{}\begin{array}[]{ll}1&\quad\mbox{if $g_{r}=s$,}\\ 0&\quad\mbox{otherwise,}\\ \end{array}$ (20)
 so the conventional mutual information of Eq. (13) is
 $I_{0}(c;g)=\log\frac{n!}{\prod_{s}n_{s}^{(g)}!}=H_{0}(g).$ (21)
-This answer is as wrong as it could be: we expect the mutual information to take the minimum value of zero, but instead it is equal to the entropy $H_{0}(g)$, which is the maximum possible value since $H_{0}(g)$ is an upper bound as we have said.
-In other words, $I_{0}(c;g)$ would have us believe that this algorithm which puts every object in its own group tells us everything there is to know about the true labeling $g$, when in fact it tells us nothing at all.
+This answer is not merely wrong; it is maximally so.
+The mutual information should take its minimum value of zero, but instead it takes the value $H_{0}(g)$, which is the maximum possible since $H_{0}(g)$ is an upper bound as we have said.
 The reason for this result is that in this case the contingency table itself uniquely defines $g$, so neglecting it puts the mutual information in error by an amount equal to the complete information cost of the ground truth.
-If we include the cost of transmitting the contingency table on the other hand, this erroneous behavior disappears.
-In this case, we can calculate the number of contingency tables exactly:
-$\Omega(n^{(c)},n^{(g)})=\frac{n!}{\prod_{s}n_{s}^{(g)}!},$ (22)
-so the reduced mutual information is
-$I(c;g)=I_{0}(c;g)-\log\frac{n!}{\prod_{s}n_{s}^{(g)}!}=0,$ (23)
-which is now the expected answer.
-A related shortcoming of the standard mutual information is that for finite $n$ a random labeling $c$ will have positive mutual information with respect to any ground truth in expectation: because mutual information is non-negative, fluctuations due to randomness in finite systems will produce non-negative values only and hence their average will in general be positive VEB10 ; Zhang15 .
+If we include the subleading terms on the other hand, this erroneous behavior disappears.
+Assuming for simplicity that the ground-truth groups are of equal size, the optimal concentration parameters are $\alpha_{g|c}=0$ and $\alpha_{g}=\infty$, and the reduced mutual information of Eq.
+(12) becomes
+$I(c;g)$ $=I_{0}(c;g)-\log\left[\frac{n!}{\prod_{r}n_{r}!}(1/q_{g})^{n}\right]-n\log q_{g}$
+$=0.$ (22)
+which is the correct answer.
+A related shortcoming of the standard mutual information is that for finite $n$ a random labeling $c$ will have positive mutual information with respect to any ground truth in expectation: because mutual information is non-negative, fluctuations due to randomness in finite systems will produce non-negative values only and hence their average will in general be positive [4, 9].
 This seems counterintuitive: we would expect the average value for a random labeling to be zero.
@@
 To do this we must first decide how a random labeling is to be defined—over what distribution of candidate labelings are we averaging?
-A popular choice is to take the uniform distribution over labelings that share the same group sizes $n^{(c)}$ as the actual candidate $c$.
+The conventional choice is to take the uniform distribution over labelings that share the same group sizes $n^{(c)}$ as the actual candidate $c$.
 This yields the adjusted mutual information of Vinh et al.
-VEB10 :
+[4]:
 $I_{A}(c;g)=I_{0}(c;g)-\langle I_{0}(c;g)\rangle_{\{c|n^{(c)}\}},$ (23)
-where the expectation is over the ensemble described.
-This gives $\langle I_{A}(c;g)\rangle_{\{c|n^{(c)}\}}=0$, so a positive value of the adjusted mutual information indicates a candidate labeling that performs above average within this ensemble.
-However, the adjusted mutual information defined in this way, and particularly the choice of ensemble, has some problems NCY20 .
-The set of labelings with given $n^{(c)}$ heavily favors contingency tables with relatively uniform entries, simply because there are many more labelings that correspond to uniform tables than to non-uniform ones.
-Real contingency tables, on the other hand, tend to be highly non-uniform, since applications of the mutual information tend to focus on labelings that are at least somewhat similar to the ground truth (which produces a non-uniform table).
-This means that the average used in the adjusted mutual information puts most of its weight on configurations that are very different from those that occur in reality, making it a poor representation of the actual average mutual information.
-A solution to this problem is to perform the average not over all labelings consistent with $n^{(c)}$ but over all contingency tables, which makes non-uniform contingency tables just as likely as uniform ones.
-As shown in NCY20 , however, adjusting the mutual information in this way simply gives us the reduced mutual information of Eq.
-(17) again.
-For this reason we favor the reduced mutual information over the adjusted mutual information: in addition to allowing for the information cost of the contingency table, it corrects sensibly for the nonzero expected value.
-It is also derived naturally from a fully information-theoretic argument, by contrast with the more ad hoc derivation of the adjusted mutual information.
+where the expectation is over the relevant ensemble.
+This measure can also be treated within the information-cost framework of the present paper, as described in [6].
+There it is shown that the subtracted term $\langle I_{0}(c;g)\rangle_{\{c|n^{(c)}\}}$ is precisely equal to the average cost of transmitting the contingency table when labelings are drawn from the uniform distribution.
+Thus, one can view the adjusted mutual information as a reduced measure, but differing from the one proposed in this paper in two ways.
+First, the correction term counts the cost of transmitting the contingency table but not the cost of transmitting the counts $n^{(g)}$, as described in Section II.
+This, however, is only a minor concern—the latter is usually much smaller than the former and its neglect introduces only a small error in most cases.
+The second difference is more serious and revolves around the choice to draw labelings uniformly from those with given $n^{(c)}$.
+As discussed in [6], this distribution heavily favors contingency tables with relatively uniform entries, simply because there are many more labelings that correspond to uniform tables than to non-uniform ones.
+Real contingency tables, on the other hand, are often highly non-uniform, since many applications of the mutual information focus on labelings that are somewhat similar to the ground truth (producing a non-uniform table).
+In such cases, the average used in the adjusted mutual information puts most of its weight on configurations that are very different from those that occur in reality, making it a poor representation of true information costs.
+The measure proposed in this paper, by contrast, deliberately allows for non-uniform tables by drawing them from a Dirichlet-multinomial distribution and we argue that this is a strong reason to favor it over the adjusted mutual information.
+In addition it is also derived naturally from a fully information-theoretic argument, by contrast with the more ad hoc derivation of the adjusted mutual information.
 Notwithstanding, in Section III we give results using both reduced and adjusted mutual information, and find fairly similar outcomes in the two cases.
@@
 In particular, some, including the most popularly used normalization, can result in biased results and should, in our opinion, be avoided.
-In its place, we propose a novel, unbiased normalized measure.
+In its place, we propose an alternative, unbiased normalized measure.
 The most popular normalized measure, commonly referred to simply as the “normalized mutual information,” uses the plain mutual information $I_{0}(c;g)$ as a base measure and normalizes it thus:
-$I_{0}^{(S)}(c;g)$ $=\frac{I_{0}(c;g)}{\frac{1}{2}[H_{0}(c)+H_{0}(g)]}$
-$=\frac{I_{0}(c;g)}{\frac{1}{2}[I_{0}(c;c)+I_{0}(g;g)]}.$ (25)
+$\text{NMI}_{0}^{(S)}(c;g)$ $=\frac{I_{0}(c;g)}{\frac{1}{2}[H_{0}(c)+H_{0}(g)]}$
+$=\frac{I_{0}(c;g)+I_{0}(g;c)}{I_{0}(c;c)+I_{0}(g;g)}.$ (24)
 This measure has a number of desirable features.
-Because of the inequalities in (13) its value falls between zero and one.
+Because of the inequalities in (16), its value falls strictly between zero and one.
 And since both the base measure and the normalization are symmetric under interchange of $c$ and $g$, the normalized measure also retains this symmetry (hence the superscript “$(S)$,” for symmetric).
@@
 Note that the network itself plays no role here—it is included only as a visual aid.
-Equation (25) is not the only normalization that achieves these goals however. Equation (13) implies that
+Equation (24) is not the only normalization that achieves these goals. Equation (16) implies that
 $I_{0}(c;g)$ $\leq\min(I_{0}(c;c),I_{0}(g;g))$ (25)
@@
 $\leq\max(I_{0}(c;c),I_{0}(g;g)),$ (27)
-which gives us three more options for the denominator of the normalized measure. The arithmetic mean in Eq. (25), however, sees the most use by far LF09 ; OLC11 ; YAT16 ; fortunato2016community .
+which gives us three more options for a symmetric denominator in the normalized measure. The arithmetic mean in Eq. (24), however, sees the most use by far [10, 11, 12, 13].
 We can extend the notion of symmetric normalization to any other base measure of mutual information $I_{X}(c;g)$ by writing
-$I_{X}^{(S)}(c;g)=\frac{I_{X}(c;g)}{\frac{1}{2}[I_{X}(c;c)+I_{X}(g;g)]}.$ (29)
+$\text{NMI}_{X}^{(S)}(c;g)=\frac{I_{X}(c;g)+I_{X}(g;c)}{I_{X}(c;c)+I_{X}(g;g)}.$ (28)
 All such measures, however, including the standard measure of Eq.
@@
 This behavior is due to the difference in entropy between the two candidate divisions, the one on the right having larger entropy than the one on the left, which increases its normalization factor and correspondingly decreases the normalized mutual information.
-We contend that in principle the unnormalized measure is more correct on this question, having a direct justification in terms of information theory.
+We argue that the unnormalized measure is more correct on this question, having a direct justification in terms of information theory.
 The purpose of the normalization is purely to map the values of the measure onto a convenient numerical interval, and should not change outcomes as it does here.
@@
 We define the asymmetric normalization of any base measure $I_{X}$ as
-$I_{X}^{(A)}(c;g)=\frac{I_{X}(c;g)}{I_{X}(g;g)}.$ (30)
+$\text{NMI}_{X}^{(A)}(c;g)=\frac{I_{X}(c;g)}{I_{X}(g;g)}.$ (29)
 This definition still gives $I_{X}^{(A)}(g;g)=1$, but now the normalization factor in the denominator has no effect on choices between candidate labelings, since it is independent of $c$.
@@
 Thus this measure also removes any ambiguity about how one should perform the normalization.
-The downside of this measure is the loss of the symmetry, which is undesirable in some applications.
-For instance, if one is comparing two candidate labelings directly to one another, rather than to a distinct ground truth, then a symmetric measure may be preferable.
-Embedding and visualization methods that employ mutual information as a similarity measure also normally demand symmetry PLC17 .
-In the most common applications of the normalized mutual information, however, where we are evaluating against a ground truth, which is an inherently asymmetric situation, the asymmetric measure makes perfect sense.
+Loosely, this asymmetrically normalized mutual information measures “how much $c$ tells us about $g$ as a fraction of all there is to know about $g$.”
+The amount of bias inherent in the symmetrically normalized measure when compared with asymmetric one can be quantified by the ratio between the two:
+${\text{NMI}_{X}^{(S)}(c;g)\over\text{NMI}_{X}^{(A)}(c;g)}$ $={[I_{X}(c;g)+I_{X}(g;c)]/[I_{X}(c;c)+I_{X}(g;g)]\over I_{X}(c;g)/I_{X}(g;g)}$
+$=\frac{1+I_{X}(g;c)/I_{X}(c;g)}{1+I_{X}(c;c)/I_{X}(g;g)}.$ (30)
+If the base measure $I_{X}$ is itself symmetric (which it is, either exactly or approximately, for all the measures we consider), then this simplifies further to
+${\text{NMI}_{X}^{(S)}(c;g)\over\text{NMI}_{X}^{(A)}(c;g)}={H(g)\over\frac{1}{2}[H(g)+H(c)]}.$ (31)
+Note how this expression results in a particular bias against complex candidate labelings $c$ that have higher entropy $H(c)$ than the ground truth.
+Traditional mutual information measures are particularly problematic when the candidate is a refinement of the ground truth, meaning the candidate groups are subsets of the ground-truth groups.
+The example given in Section II.1, in which the candidate labeling puts every node in a group on its own, is an extreme instance of this.
+Equation (31) shows that, to some extent, the traditional symmetric normalization corrects this issue: a candidate $c$ that takes the form of a refinement of $g$ will have $H(c)>H(g)$ and hence the normalized mutual information will be penalized.
+We argue, however, that this is not the best way to address the problem and that the correct approach is instead to use the reduced mutual information.
+The reduced mutual information also penalizes such cases, but it does so in a more principled manner that directly addresses the root cause of the problem, rather than merely penalizing complex candidate labelings in an ad hoc manner as a side-effect of normalization.
+We will see examples of this in Section III, where we apply both measures to community detection in networks and find that, although the symmetrically normalized traditional measure is biased, it is in some cases fortuitously biased in the right direction, although it is still problematic in some others.
+An obvious downside of asymmetric normalization is the loss of the symmetry in the final measure.
+In the most common applications of normalized mutual information, where labelings are evaluated against a ground truth, an inherently asymmetric situation, the asymmetric measure makes perfect sense, but in other cases the lack of symmetry can be undesirable.
+Embedding and visualization methods that employ mutual information as a similarity measure, for example, normally demand symmetry [14].
+And in cases where one is comparing two candidate labelings directly to one another, rather than to a separate ground truth, then a symmetric measure may be preferable.
+Even in this case, however, the asymmetric measure may sometimes be the better choice, as discussed in the introduction.
+For instance, when one labeling $c_{1}$ is a refinement of the other $c_{2}$, the information content is inherently asymmetric: $c_{1}$ says more about $c_{2}$ than $c_{2}$ does about $c_{1}$.
+An explicit example of this type of asymmetry is shown in Figure 3, where we consider two labelings of 27 objects.
+The left labeling $c_{1}$ is a detailed partition of the objects into nine small groups while the right labeling $c_{2}$ is a coarser partition into only three groups, each of which is an amalgamation of three of the smaller groups in $c_{1}$.
+Because of this nested relationship, it is relatively easy to transmit $c_{2}$ given knowledge of $c_{1}$ but more difficult to do the reverse.
+This imbalance is reflected in the asymmetrically normalized mutual information values in each direction (top and bottom arrows in the figure), but absent from the symmetric version (middle row).
 Combining the benefits of asymmetric normalization and the reduced mutual information, we advocate in favor of the asymmetrically-normalized reduced mutual information defined by
-$I^{(A)}(c;g)=\frac{I(c;g)}{I(g;g)},$ (31)
-where
-$I(c;g)=I_{0}(c;g)-\log\Omega(n^{(c)},n^{(g)}),$ (32)
-with $\Omega(n^{(c)},n^{(g)})$ approximated as in Eq.
+$\text{NMI}^{(A)}(c;g)=\frac{I(c;g)}{I(g;g)},$ (32)
+where the mutual information $I(c;g)$ is quantified as in Eq.
 (12).
-This measure correctly accounts for the information contained in the contingency table, returns value 0 in expectation when $c$ is uncorrelated with the ground truth and value 1 when $c=g$, and always favors the same labeling as the unnormalized measure.
-One potential issue is that at present we are unable to prove a bound $I(c;g)\leq I(g;g)$ analogous to (13), which would ensure that $I^{(A)}(c;g)\leq 1$ always.
-There is strong numerical and other evidence that such a bound does apply—see Appendix B—but proving it is an open problem.
+This measure correctly accounts for the information contained in the contingency table, returns 0 in expectation when $c$ is uncorrelated with the ground truth and 1 when $c=g$, and always favors the same labeling as the unnormalized measure.
+Figure 3: Values of the various normalized mutual information measures we consider, between two example partitions of the same set of 27 objects.
+The left partition $c_{1}$ divides the objects into nine groups of three objects each and is a refinement of the right partition $c_{2}$, which divides them into only three groups, aggregating those on the left.
+The arrows indicate the direction of the comparison and the accompanying notations give the values of the mutual information measures.
+All three asymmetrically normalized measures capture the intuition that the partition $c_{1}$ tells us more about $c_{2}$ than $c_{2}$ tells us about $c_{1}$.
 III Example application:
 Community detection
-As an example of the performance of our favored measure of mutual information and competing measures, we describe in this section the results of an extensive series of tests in which these measures are used to score the output of various algorithms for network community detection.
-In these tests we use the popular Lancichinetti-Fortunato-Radicchi (LFR) graph model LFR08 to generate random test networks with known community structure and realistic distributions of node degrees and group sizes, and then attempt to recover that structure using a variety of standard community detection algorithms, quantifying the accuracy of the recovery with six different measures: symmetrically normalized versions of the traditional mutual information, the adjusted mutual information, and the reduced mutual information, and asymmetrically normalized versions of the same.
-A number of studies have been performed in the past to test the efficacy of community detection algorithms on LFR benchmark networks LF09 ; OLC11 ; YAT16 ; fortunato2016community , but using only the symmetric non-reduced mutual information as a similarity measure.
+As an example of the performance of the various measures discussed above, we describe in this section the results of an extensive series of tests in which the measures are used to score the output of algorithms for network community detection.
+In these tests we use the popular Lancichinetti-Fortunato-Radicchi (LFR) graph model [15] to generate random test networks with known community structure and realistic distributions of node degrees and group sizes, and then attempt to recover that structure using a variety of standard community detection algorithms, quantifying the accuracy of the recovery with six different measures: symmetrically normalized versions of the traditional mutual information, the adjusted mutual information, and the reduced mutual information, and asymmetrically normalized versions of the same.
+A number of studies have been performed in the past to test the efficacy of community detection algorithms on LFR benchmark networks [10, 11, 12, 13], but using only the symmetrically normalized, non-reduced mutual information as a similarity measure.
 Our results indicate that this measure can produce biased outcomes and we recommend the asymmetric reduced mutual information instead.
@@
 (We give further details on the LFR generative process in Appendix A.) We find that the distributions of degrees and community sizes do not significantly impact the relative performance of the various algorithms tested and that performance differences are driven primarily by the size $n$ of the networks and the mixing parameter $\mu$ that controls the ratio of within/between-group connections, so our tests focus on performance as a function of these variables.
+Figure 4: The colors in each panel in the center of this figure indicate which of three community detection algorithms (InfoMap (red) and modularity maximization with $\gamma=1$ (green) and $\gamma=10$ (blue)—see key on right) are best able to find the known communities in a large set of LFR benchmark networks, according to the six mutual information measures we consider.
+Mixtures of red, green, and blue denote the proportions of test cases in which each algorithm performs best.
+Regions in gray indicate parameter values for which no algorithm achieved a positive mutual information score.
 We perform community detection using six well-known algorithms as follows.
-(We use the implementations found in the igraph library CN06 , except for the inference method, for which we use the graph-tool library P14 .)
+(We use the implementations found in the igraph library [16], except for the inference method, for which we use the graph-tool library [17].)
   1. 1.
-    InfoMap: InfoMap is an information theoretic approach that defines a compression algorithm for encoding a random walk on a network based on which communities the walk passes through RB08 .
+    InfoMap: InfoMap is an information theoretic approach that defines a compression algorithm for encoding a random walk on a network, based on which communities the walk passes through [18].
 Different community labelings give rise to more or less efficient compression, as quantified by the so-called map equation, and the labeling with the highest efficiency is considered the best community division.
@@
     Modularity maximization: Modularity is a quality function for network labelings equal to the fraction of edges within communities minus the expected such fraction if edge positions are randomized while preserving the node degrees.
-Modularity maximization aims to optimize the modularity over possible labelings and return the labeling with the highest modularity found.
-Modularity can be maximized using a number of heuristic methods, of which the most popular are agglomerative methods such as the Louvain and Leiden algorithms BGLL08 ; TWV19 , spectral methods Newman06b , and simulated annealing GSA04 ; MAD05 ; RB06a .
+The labeling with the highest modularity is considered the best community division.
+Modularity can be maximized using a number of heuristic methods, of which the most popular are agglomerative methods such as the Louvain and Leiden algorithms [19, 20], spectral methods [21], and simulated annealing [22, 23, 24].
 In our tests we use simulated annealing where computationally feasible and the Leiden algorithm otherwise, these approaches giving the most consistent maximization of the modularity.
   3. 3.
-    Modularity with enhanced resolution: Standard modularity maximization is known to suffer from a “resolution limit”—it cannot detect communities smaller than a certain threshold size FB07 .
-This can be remedied using a variant of modularity that includes a resolution parameter $\gamma$ such that higher values of $\gamma$ cause the algorithm to prefer smaller communities RB06a .
+    Modularity with a resolution parameter: Standard modularity maximization is known to suffer from a “resolution limit”—it cannot detect communities smaller than a certain threshold size [25].
+This can be remedied using a variant of modularity that includes a resolution parameter $\gamma$ such that higher values of $\gamma$ cause the algorithm to prefer smaller communities [24].
 Standard modularity maximization corresponds to $\gamma=1$, but for comparison we also conduct tests with $\gamma=10$ using the Leiden algorithm.
@@
     Statistical inference: Another popular approach to community detection makes use of model fitting and statistical inference.
-In this context the most commonly fitted model is the degree-corrected stochastic block model KN11a , which can be fitted using Bayesian methods to find the best community division Peixoto17 .
+In this context the most commonly fitted model is the degree-corrected stochastic block model [26], which can be fitted using Bayesian methods to find the best community division [27].
   5. 5.
-    Walktrap: Walktrap is an agglomerative algorithm in which initially separate nodes are iteratively combined into progressively larger communities in order from strongest to weakest connections, where strength is quantified in terms of the time for a random walk to reach one node from another PL05 .
+    Walktrap: Walktrap is an agglomerative algorithm in which initially separate nodes are iteratively combined into progressively larger communities in order from strongest to weakest connections, where strength is quantified in terms of the time for a random walk to reach one node from another [28].
   6. 6.
-    Labelprop: The label propagation or “labelprop” algorithm initially places every node in its own community then iteratively updates the labels of randomly chosen nodes by majority vote among their network neighbors, breaking ties at random RRS07 .
+    Labelprop: The label propagation or “labelprop” algorithm initially places every node in its own community then iteratively updates the labels of randomly chosen nodes by majority vote among their network neighbors, breaking ties at random [29].
 As described in Section III.2, all of these algorithms perform reasonably well, but, as we will see, the best performers in the context of the current tests are InfoMap and the two variants of modularity maximization.
-Figure 3: The colors in each panel in the center of this figure indicate which of three community detection algorithms (InfoMap (red) and modularity maximization with $\gamma=1$ (green) and $\gamma=10$ (blue)—see key on right) are best able to find the known communities in a large set of LFR benchmark networks, according to the six mutual information measures we consider.
-Mixtures of red, green, and blue denote the proportions of test cases in which each algorithm performs best.
-Regions in gray indicate parameter values for which no algorithm achieved a positive mutual information score.
 III.1 Comparison between variants of the
@@
 In this set of tests we limited ourselves, for the sake of clarity, to the top three community detection algorithms—InfoMap and the two variants of modularity maximization—and measured which of the three returned the best results according to each of our six mutual information measures, as a function of network size $n$ and the mixing parameter $\mu$.
-Each point in each of the six panels in the figure is color-coded with some mix of red, green, and blue to indicate in what fraction of cases each of the algorithms performs best according to the relevant measure and, as we can see, the results vary significantly among the measures.
+Each point in each of the six panels in the figure is color-coded with some mix of red, green, and blue to indicate in what fraction of cases each of the algorithms performs best according to each of the six measures and, as we can see, the results vary significantly among measures.
 An experimenter trying to choose the best algorithm would come to substantially different conclusions depending on which measure they use.
 One consistent feature of all six mutual information measures is the large red area in each panel of Fig.
-3, which represents the region in which InfoMap performs best.
+4, which represents the region in which InfoMap performs best.
 Regardless of the measure used, InfoMap is the best performer on networks with low mixing parameter (i.e., strong community structure) and relatively large network size.
@@
 An extreme example of the disagreement between measures can be seen in the lower right corner of each plot in Fig.
-3, where most measures return a negative value, indicating that community detection has failed altogether.
-This is expected: for very weak community structure all detection algorithms are expected to show a “detectability threshold” beyond which they are unable to identify any communities DKMZ11a ; Massoulie14 ; MNS15 .
+4, where most measures return a negative value, indicating that community detection has failed altogether.
+This is expected: for very weak community structure all detection algorithms are expected to show a “detectability threshold” beyond which they are unable to identify any communities [30, 31, 32].
 The standard normalized mutual information, however, claims to find community structure in this regime using the $\gamma=10$ version of modularity maximization.
@@
 The bottom left panel in Fig.
-3 shows results for the asymmetrically normalized version of the traditional mutual information, which, perhaps surprisingly, gives even worse results than the symmetric version, with hardly any region in which the $\gamma=1$ version of modularity maximization outperforms the $\gamma=10$ version.
-The reason for this is interesting: the symmetric normalization of the traditional measure actually acts as an ad hoc correction against the errors introduced by neglecting the information content of the contingency table and hence reduces the amount by which the symmetric measure incorrectly rewards labelings with too many communities.
+4 shows results for the asymmetrically normalized version of the traditional mutual information, which gives even worse results than the symmetric version, with hardly any region in which the $\gamma=1$ version of modularity maximization outperforms the $\gamma=10$ version.
+This behavior arises for the reasons discussed in Section II.2: the bias inherent in the symmetric normalization fortuitously acts to partially correct the errors introduced by neglecting the information content of the contingency table.
 The asymmetric normalization eliminates this correction and hence performs more poorly.
-The correct solution to this problem, however, is not to use a symmetric normalization, which can bias outcomes in other ways, but rather to adopt a reduced mutual information measure.
+The correct solution to this problem, however, is not to use a symmetric normalization, which can bias outcomes in other ways as we have seen, but rather to adopt a reduced mutual information measure.
 Finally, comparing the middle and right-hand columns of Fig.
-3, we see that the results for the adjusted and reduced mutual information measures are quite similar in these tests, although there are some differences.
+4, we see that the results for the adjusted and reduced mutual information measures are quite similar in these tests, although there are some differences.
 In particular, the adjusted measure appears to find more significant structure for higher mixing than the reduced measure.
 This occurs because, as discussed in Section II.1 and Ref.
-NCY20 , the adjusted measure encodes the contingency table in a way that is optimized for more uniform tables than the reduced measure, and thus penalizes uniform tables less severely, leading to overestimates of the mutual information in the regime where detection fails.
+[6], the adjusted measure encodes the contingency table in a way that is optimized for more uniform tables than the reduced measure, and thus penalizes uniform tables less severely, leading to overestimates of the mutual information in the regime where detection fails—in this regime the candidate and ground-truth labelings are uncorrelated which results in a uniform contingency table.
 This provides further evidence in favor of using a reduced mutual information measure.
-III.2 Comparison between community detection algorithms
+III.2 Comparison between
+community detection algorithms
 Settling on the asymmetrically normalized reduced mutual information as our preferred measure of similarity, we now ask which community detection algorithm or algorithms perform best according to this measure?
@@
 Figure 6 shows results for all six algorithms listed in Section III.
-Examining the figure we see that in general the best-performing methods are InfoMap, traditional modularity maximization with $\gamma=1$, and the inference method using the degree-corrected stochastic block model.
+Examining the figure, we see that in general the best-performing methods are InfoMap, traditional modularity maximization with $\gamma=1$, and the inference method using the degree-corrected stochastic block model.
 Among the algorithms considered, InfoMap achieves the highest mutual information scores for lower values of the mixing parameter $\mu$ in the LFR model, but fails abruptly as $\mu$ increases, so that beyond a fairly sharp cutoff around $\mu=0.5$ other algorithms do better.
-InfoMap’s specific failure mode is that it places all nodes in a single community and this behavior can be used as a simple indicator of the failure regime.
+As noted by previous authors [33], InfoMap’s specific failure mode is that it places all nodes in a single community and this behavior can be used as a simple indicator of the failure regime.
 In this regime one must use another algorithm.
-Either modularity maximization or inference using the degree-corrected stochastic block model are reasonable options, but modularity has a slight edge, apart from a thin band of intermediate $\mu$ values which, in the interests of simplicity, we choose to ignore.
+Either modularity maximization or inference using the degree-corrected stochastic block model are reasonable options, but modularity has a slight edge, except in a thin band of intermediate $\mu$ values which, in the interests of simplicity, we choose to ignore.
 (We discuss some caveats regarding the relationship between the degree-corrected stochastic block model and the LFR benchmark in Appendix A.)
@@
 Acknowledgements.
-This work was supported in part by the US National Science Foundation under grant DMS–2005899 and by computational resources provided by the Advanced Research Computing initiative at the University of Michigan.
+The authors thank Samin Aref for useful comments and feedback.
+This work was supported in part by the US National Science Foundation under grants DMS–2005899 and DMS–2404617, and by computational resources provided by the Advanced Research Computing initiative at the University of Michigan.
 Appendix A Benchmarking
@@
 Figure 7 shows the performance of the same six community detection methods as in Fig.
-5, but measured using the standard, symmetrically normalized, non-reduced mutual information $I_{0}^{(S)}(c;g)$.
+6, but measured using the standard, symmetrically normalized, non-reduced mutual information $I_{0}^{(S)}(c;g)$.
 By this measure many of the methods appear to perform implausibly well, far beyond the detectability threshold visible in Fig.
@@
 A.2 LFR network generation
-The LFR networks we use for benchmarking are generated using the procedure described in LFR08 , which is as follows.
+The LFR networks we use for benchmarking are generated using the procedure described in [15], which is as follows.
   1. 1.
@@
 For small $\mu$ there are many more edges within communities than between them, which makes the communities relatively easy to detect.
-But as $\mu$ gets larger there are more edges between communities and detection becomes more difficult in a manner reminiscent of the detectability threshold in the standard stochastic block model DKMZ11a ; DKMZ11b .
+But as $\mu$ gets larger there are more edges between communities and detection becomes more difficult in a manner reminiscent of the detectability threshold in the standard stochastic block model [30, 34].
 In our studies we use values of $\mu$ in the range $[0.2,0.8]$.
@@
     Draw a degree sequence from a power-law distribution with exponent $\tau_{1}$.
-Many networks have power-law degree distributions, typically with exponents between 2 and 3 Caldarelli07 , and the LFR model exclusively uses power-law distributions.
+Many networks have power-law degree distributions, typically with exponents between 2 and 3 [35], and the LFR model exclusively uses power-law distributions.
 We use $\tau_{1}=2.5$, with average degree $\langle k\rangle=20$ and maximum degree (which depends on graph size) $k_{\text{max}}=n/10$.
@@
     Draw a set of community sizes from a power-law distribution with exponent $\tau_{2}$.
-Many networks are also found to have community sizes that approximately follow a power law, with typical exponents in the range from 1 to 2 Guimera03 ; CNM04 ; PDFV05 ; LFR08 .
+Many networks are also found to have community sizes that approximately follow a power law, with typical exponents in the range from 1 to 2 [36, 37, 38, 15].
 We use $\tau_{2}=1.5$ with a minimum community size of $s_{\text{min}}=20$ in all cases, while the maximum community size is set to $s_{\text{max}}=\max\left\{n/10,100\right\}$.
@@
     Rewire the edges attached to each node while preserving the node degrees, until the fraction of edges running between nodes in different communities is approximately $\mu$.
-The parameter values above are similar to those used for instance in YAT16 .
+The parameter values above are similar to those used for instance in [12].
 As in that study, we find that algorithm performance is dictated primarily by the parameters $n$ and $\mu$, so it is these parameters that are varied our summary figures.
-The LFR model is similar to a special case of the degree-corrected stochastic block model (DCSBM) KN11a , and hence one might expect that inference-based community detection methods employing the latter model would perform optimally on LFR networks.
+The LFR model is similar to a special case of the degree-corrected stochastic block model (DCSBM) [26], and hence one might expect that inference-based community detection methods employing the latter model would perform optimally on LFR networks.
 Specifically, in the limit of an infinite number of sampled networks and perfect optimization of each community detection method, the final performance measure for any algorithm is given by the expectation value of the similarity $M(g,h[A])$ between the ground truth LFR partition $g$ and the partition $h[A]$ of the LFR network $A$ inferred using the algorithm, where the expectation is taken over the ensemble $P(A,g|\theta)$ of LFR networks and partitions $A,g$ generated using parameters $\theta$ (meaning $\mu$, $\tau_{1}$, $\tau_{2}$, etc.).
@@
 In other words, the optimal algorithm simply performs maximum a posteriori estimation under the model from which the network was generated.
-There is no explicit formula for the posterior probability under the LFR model, but to the extent that it is a special case of the DCSBM, we might expect the DCSBM (with appropriate priors) to give optimal results peixoto2021revealing .
+There is no explicit formula for the posterior probability under the LFR model, but to the extent that it is a special case of the DCSBM, we might expect the DCSBM (with appropriate priors) to give optimal results [39].
 The LFR model, however, is not precisely a special case of the DCSBM.
-In particular, the DCSBM normally assumes a uniform distribution over community sizes, where the LRF model assumes a power law.
+In particular, the DCSBM normally assumes a uniform distribution over community sizes, where the LFR model assumes a power law.
 Moreover, we are not using the crude all-or-nothing error function: our entire purpose in this paper is to develop mutual information measures that aggregrate and weigh different modes of error in a sensible fashion.
 These differences, it appears, are enough to ensure that the DCSBM does not perform the best in our testing.
-Regardless, we emphasize that our use of the LFR benchmark in our analysis is simply for consistency with previous studies of network community detection methods LFR08 ; LF09 ; YAT16 .
+Regardless, we emphasize that our use of the LFR benchmark in our analysis is simply for consistency with previous studies of network community detection methods [15, 10, 12].
 The justification for our proposed similarity measure, on the other hand, is chiefly its theoretical merit over the conventional (symmetric, non-reduced) normalized mutual information, and is independent of the use of the LFR (or any other) benchmark.
-Appendix B Upper bound on the
-reduced mutual information
-In this appendix, we provide evidence for the inequality
-$I(c;g)\leq I(g;g),$ (34)
-with the exact equality applying if and only if the labelings $c$ and $g$ are identical up to a permutation of the label names.
-If true, this inequality would imply that the reduced normalized measure $I^{(A)}(c;g)$ was bounded above by 1, as is the unreduced version $I_{0}(c;g)$.
-We do not at present know of a general proof of (34), but here we present some weaker theoretical results along with strong numerical evidence of its correctness.
-First, we note that the only possible violations of (34) would be for small $n$.
-As $n$ becomes large, the second (reduction) term in Eq.
-(9) scales as $\log n$, a subleading correction, and hence is negligible for large $n$ compared to the leading mutual information itself which is of order $n$, and the mutual information alone satisfies (34).
-We have exhaustively checked the following cases with $n$ up to 200 and find (34) to be satisfied in all of them:
-$q_{c}=2,\ q_{g}=2$ for $n\leq 200$,
-$q_{c}=3,\ q_{g}=2$ for $n\leq 50$,
-$q_{c}=4,\ q_{g}=2$ for $n\leq 30$,
-$q_{c}=3,\ q_{g}=3$ for $n\leq 30$,
-$q_{c}=4,\ q_{g}=3$ for $n\leq 20$.
-We also observe zero violations of (9) in any of the (tens of thousands of) randomly generated examples which went into the creation of Fig. 5.
-On the theoretical side, (34) implies that $c=g$ is a global maximum of $I(c;g)$, which we are unable to prove, but we can prove the weaker result that it is a local maximum under the mislabeling of any single node, when the reduction term is calculated using Eq.
-(18).
-To do this, we first assume, without loss of generality, that a single node with ground truth community label 1 is mislabeled in the candidate partition $c$ as being in community 2.
-This implies that $n_{1}^{(c)}=n_{1}^{(g)}-1$, $n_{2}^{(c)}=n_{2}^{(g)}+1$, and $n_{r}^{(c)}=n_{r}^{(g)}$ for all $r=3\ldots q_{g}$.
-The contingency table will then be nearly diagonal, with $(n_{1}^{(g)}-1,n_{2}^{(g)},\ldots)$ on the diagonal and a single nonzero off-diagonal element $n_{12}^{(cg)}=1$.
-The decrease in the standard mutual information is then given by
-$I_{0}(g;g)-I_{0}(g;c)=\log\frac{n!}{\prod_{s}n_{s}^{(g)}!}$
-$\quad{}-\log\frac{n!1!(n_{1}^{(g)}-1)!\prod_{s=2}^{q_{g}}n_{s}^{(g)}!}{\left(\prod_{s}n_{s}^{(g)}!\right)\left(\prod_{s=3}^{q_{g}}n_{s}^{(g)}!\right)(n_{1}^{(g)}-1)!(n_{2}^{(g)}+1)!}$
-$=\log(n_{2}^{(g)}+1).$ (35)
-At the same time, the reduction term $\log\Omega$ also decreases, by an amount that can be estimated from Eq. (18) thus:
-$\log$ $\Omega(n^{(g)},n^{(g)})-\log\Omega(n^{(g)},n^{(c)})=\log\prod_{r=1}^{q_{g}}\frac{\binom{n_{r}^{(g)}+\alpha-1}{\alpha-1}}{\binom{n_{r}^{(c)}+\alpha-1}{\alpha-1}}$
-$=\log\frac{(n_{1}^{(g)}+\alpha+1)(n_{2}^{(g)}+1)}{n_{1}^{(g)}(n_{2}^{(g)}+\alpha)}.$ (36)
-Combining Eqs. (35) and (36), we now find that the change in the RMI incurred from the mislabeling of a single node is
-$I(g;g)-I(g;c)=\log\frac{n_{1}^{(g)}(n_{2}^{(g)}+\alpha)}{n_{1}^{(g)}+\alpha-1}>0,$ (37)
-where the inequality holds so long as $n_{1}^{(g)},n_{2}^{(g)},\alpha\geq 1$.
-Thus, under the estimate of $\Omega(n^{(g)},n^{(c)})$ given by Eq.
-(18), $g=c$ is a local maximum of the reduced mutual information.

### v3

@@
 As a practical demonstration of the importance of using an unbiased measure, we perform extensive numerical tests on a basket of popular algorithms for network community detection and show that one’s conclusions about which algorithm is best are significantly affected by the biases in the traditional mutual information.
-I Introduction
-A common task in data analysis is the comparison of two different labelings of the same set of objects.
+1 Introduction
+A common task in data analysis is the comparison of two different labelings of a set of objects.
 How well do demographics predict political affiliation?
 How accurately do blood tests predict clinical outcomes?
-The information theoretic measure known as mutual information is commonly applied to answer questions like these, in which an experimental or computational estimate of some kind is compared against a “ground truth.”
-Mutual information [1] is a measure of how efficiently we can describe one labeling of the set of objects if we already know another labeling.
+These are questions about how similar one labeling of individuals/items/results is to another and they are commonly tackled using the information theoretic measure known as mutual information, in which an experimental or computational measure of some kind is compared against a ground truth.
+Mutual information works by asking of how efficiently we can describe one labeling if we already know the other [1].
 Specifically, it measures how much less information it takes to communicate the first labeling if we know the second versus if we do not.
@@
 Mutual information has a number of appealing properties as a tool for comparing labelings.
-It is invariant under permutations of the labels, so that, for example, two divisions of a network into communities that differ only in how the communities are labeled will be correctly identified as being the same.
+It is invariant under permutations of the labels, so that, for example, two divisions of a network into communities that differ only in the numbering of the communities will be correctly identified as being the same.
 It also returns sensible results in cases where the number of distinct label values is not the same in the two labelings.
@@
 First, it has a bias towards labelings with too many distinct label values.
-For instance, a community detection algorithm that routinely divides networks into significantly more groups than are present in the ground truth can nonetheless achieve high mutual information scores.
+For instance, a community detection algorithm that incorrectly divides networks into significantly more groups than are present in the ground truth can nonetheless achieve a high mutual information score.
 A number of approaches for correcting this flaw have been proposed.
-One can apply direct penalties for incorrect numbers of groups [3] or subtract correction terms based on the average value of the mutual information over some ensemble of candidate labelings [4, 5] or on the statistics of the contingency table [6, 7].
+One can apply direct penalties for incorrect numbers of groups [3] or subtract correction terms based on the average mutual information over some ensemble of candidate labelings [4, 5] or on the statistics of the contingency table [6, 7].
 For reasons we discuss shortly, we favor the latter approach, which leads to the measure known as the reduced mutual information.
@@
 As we demonstrate in this paper, however, these normalizations introduce biases into the results by comparison with the unnormalized measure, because the normalization factor depends on the candidate labeling as well as the ground truth.
-This effect can be large enough to change scientific conclusions, and we provide examples of this phenomenon.
-In order to avoid this latter bias, while still retaining the interpretability of a normalized mutual information measure, in this paper we propose normalizing by the entropy of the ground-truth labeling alone.
+This effect can be large enough to change scientific outcomes, and we provide examples of this phenomenon.
+In order to avoid this latter bias, while still retaining the interpretability of a normalized mutual information measure, we favor normalizing by the entropy of the ground-truth labeling alone.
 This removes the source of bias but introduces an asymmetry in the normalization.
-At first sight this asymmetry might seem undesirable, and previous authors have gone to some lengths to avoid it.
-Here, however, we argue that it is not only justified but even desirable, for several reasons.
+At first sight this asymmetry may seem undesirable, and previous authors have gone to some lengths to avoid it.
+Here, however, we argue that it is not only justified but actually desirable, for several reasons.
 First, many of the classification problems we consider are unaffected by the asymmetry, since they involve the comparison of one or more candidate labelings against a single, unchanging ground truth.
-Moreover, by contrast with the multitude of possible symmetric normalizations, the asymmetric measure we propose is the unique unbiased way to normalize the mutual information such that a perfect match receives a score of 1.
-But more broadly, one can argue that an asymmetric measure can be more informative than the conventional symmetric one.
-Consider for instance the common situation where one labeling is a fine-grained version of the other.
+Moreover, by contrast with the multitude of possible symmetric normalizations, the asymmetric measure we propose is the unique way to normalize the mutual information without introducing biases due to the normalization itself.
+More broadly, the asymmetric measure is more informative than the conventional symmetric one.
+Consider for instance the common situation where the groups or communities in one labeling are a subdivision of those in the other.
 We might for example label individuals by the country they live in on the one hand and by the town or city on the other.
-Then the detailed labeling tells us everything there is to know about the coarse-grained one, but the reverse is not true: telling you the city fixes the country, but not vice versa.
-Thus one could argue that the mutual information between the two should be asymmetric even before normalization—labeling A contains more information about labeling B than B does about A—and this type of asymmetry will be a key feature of the measures we study.
+Then the more detailed labeling tells us everything there is to know about the coarser one, but the reverse is not true.
+Telling you the city fixes the country, but not vice versa.
+Thus one could argue that the mutual information between the two eshould be asymmetric, and this type of asymmetry will be a key feature of the measures we study.
 Both drawbacks of the standard mutual information—bias towards too many groups and dependence of the normalization on the candidate labeling—can be addressed simultaneously by using an asymmetric normalized reduced mutual information as defined in this paper.
@@
 Within this framework, we find that conclusions about which algorithms perform best are significantly impacted by the choice of mutual information measure, and specifically that traditional measures erroneously favor algorithms that find too many communities, but our proposed measure does not.
-Code implementing our measure may be found at https://github.com/maxjerdee/reduced_mutual_information.
-II Mutual information
+Code implementing our measure can be found at https://github.com/maxjerdee/clustering-mi.
+2 Results
 Mutual information can be thought of in terms of the amount of information it takes to transmit a labeling from one person to another.
-In general, if we have a set of $N$ possible values for a quantity, such as possible labelings of a set of objects, then one can communicate a single one of them to a receiver by assigning each value a unique binary string and then transmitting the appropriate string.
-The minimum number of bits needed to do this is
-$H=\lceil\log_{2}N\rceil\simeq\log_{2}N\propto\ln N.$ (1)
-Conventionally one uses base-2 logarithms in this context, which gives $H$ units of bits, but we use natural logarithms in this paper, since they are more convenient for our purposes.
-The only difference is an overall factor of $\ln 2$, which will have no effect on our conclusions.
-Henceforth, all logarithms should be read as natural logs.
-Suppose that we have a labeling or division of $n$ objects into $q$ groups, which we represent as a vector of $n$ integer elements, each with value in the range $1\ldots q$.
-We assume there to be a ground-truth labeling, which we denote $g$, and a candidate labeling for comparison $c$, generated for instance by some sort of algorithm.
-The mutual information $I(c;g)$ between the two is the amount of information that can be saved when transmitting the truth $g$ if the receiver already knows the candidate $c$.
+We represent a labeling or division of $n$ objects into $q$ groups as a vector of $n$ integer elements, each with value in the range $1\ldots q$.
+We assume there to be a ground-truth labeling, which we denote $g$, with $q_{g}$ groups, and a candidate labeling for comparison $c$ with $q_{c}$ groups, generated for instance by some sort of algorithm.
+The mutual information $I(c;g)$ between the two is the amount of information saved when transmitting the truth $g$ if the receiver already knows the candidate $c$.
 We can write this information as the total entropy of $g$ minus the conditional entropy of $g$ given $c$:
-$I(c;g)=H(g)-H(g|c).$ (2)
-Loosely, we say that $I(c;g)$ measures “how much $c$ tells us about $g$.”
-We now write down expressions for both entropies in (2), but there are some subtleties involved in doing so, because the expressions depend on how the transmitted information is encoded and also because achieving correct results requires us to carefully retain some terms that are often neglected.
-Our presentation follows that of [7].
-First consider $H(g)$, which represents the information required to transmit the ground truth on its own.
-The transmission process has three steps.
-We first transmit the number of groups $q_{g}$ in the labeling, followed by a vector $n^{(g)}$ with $q_{g}$ elements $n^{(g)}_{s}$ equal to the number of objects in each group.
-Third, we transmit our final labeling $g$.
-If we apply Eq.
-(1) to calculate the cost of the first step, since the maximum possible value of $q_{g}$ is $n$, transmitting any particular value requires information $H(q_{g})=\log n$.
-Likewise for the third step, given $n^{(g)}$, there are $n!/\prod_{s}n_{s}^{(g)}!$ labelings $g$ that have the correct group sizes, so the amount of information needed to identify a particular labeling uniquely is
-$H(g|n^{(g)})=\log\frac{n!}{\prod_{s}n_{s}^{(g)}!}.$ (3)
-The middle step, transmission of the group sizes, is more delicate: here some optimization is possible that allows us to achieve a better measure of information content.
-Generally when transmitting a sequence of $n$ values with unequal probabilities such that value $r$ occurs $n_{r}$ times, the information cost is given by the generalization of Eq.
-(3) thus:
-$\log{n!\over\prod_{r}n_{r}!}$ $\simeq n\log n-n-\sum_{r}(n_{r}\log n_{r}-n_{r})$
-$=-\sum_{r}n_{r}\log p_{r},$ (4)
-where $p_{r}=n_{r}/n$ is the probability of value $r$ and we have approximated the factorials using Stirling’s formula.
-Equation 4 tells us that the information cost to transmit the value $r$ is simply $-\log p_{r}$.
-If the $N$ possible values are equally likely to appear and $p_{r}=1/N$, we recover the information cost of Eq.
-(1), but if they are not we can take advantage of that fact to create more efficient encodings.
-The transmission of $n^{(g)}$ benefits from such an efficient encoding.
-We do not in general know the distribution of the elements of $n^{(g)}$, but we know they that may be—and often are—heterogeneous.
-As shown in [7], an effective approach is to assume a Dirichlet-multinomial distribution with concentration parameter $\alpha_{g}>0$
-$P(n^{(g)}|q_{g},\alpha_{g})=$ (5)
-$\qquad\binom{n+q_{g}\alpha_{g}-1}{q_{g}\alpha_{g}-1}^{-1}\prod_{r=1}^{q_{g}}\binom{n_{r}^{(g)}+\alpha_{g}-1}{\alpha_{g}-1},$
-where for non-integer $\alpha_{g}$ we generalize the binomial coefficient in the obvious way:
-$\Bigl{(}{n\atop k}\Bigr{)}={\frac{\Gamma(n+1)}{\Gamma(k+1)\Gamma(n-k+1)}},$ (6)
-with $\Gamma(x)$ being the standard gamma function.
-If $\alpha_{g}=1$, the Dirichlet-multinomial distribution is uniform over all $q_{g}$-vectors $n^{(g)}$ of non-negative integers that sum to $n$.
-If $10<\alpha_{g}<$ 1 the distribution favors vectors with more extreme elements, while for $\alpha_{g}>1$ it favors vectors with more homogeneous elements.
-With this choice the resulting information cost to transmit $n^{(g)}$ is
-$H($ $n^{(g)}|q_{g},\alpha_{g})=-\log P(n^{(g)}|q_{g},\alpha_{g})$ (7)
-$=\log\binom{n+q_{g}\alpha_{g}-1}{q_{g}\alpha_{g}-1}-\sum_{r=1}^{q_{g}}\log\binom{n_{r}^{(g)}+\alpha_{g}-1}{\alpha_{g}-1}.$
-The optimal encoding for transmitting $n^{(g)}$ within the Dirichlet-multinomial family is given by the minimum of this expression with respect to $\alpha_{g}$, which is also equivalent to simply maximizing $P(n^{(g)}|\alpha_{g})$, i.e., to finding the maximum-likelihood value of $\alpha_{g}$.
-Note that we omit the cost of transmitting the parameter $\alpha_{g}$ itself since it will ultimately cancel out of the expression for the mutual information as described in [7].
-The total cost to transmit $g$ is now equal to the sum of the individual costs of the three steps, thus:
-$H(g)$ $=H(q_{g})+H(n^{(g)}|q_{g},\alpha_{g})+H(g|n^{(g)})$
-$=\log n+\log\frac{n!}{\prod_{s}n_{s}^{(g)}!}$ (8)
-$+\log\binom{n+q_{g}\alpha_{g}-1}{q_{g}\alpha_{g}-1}-\sum_{r=1}^{q_{g}}\log\binom{n_{r}^{(g)}+\alpha_{g}-1}{\alpha_{g}-1}$
-This three-part encoding scheme is not the only one that can be applied to this problem, but it is an efficient one in the common case of a relatively small number of groups $q_{g}\ll n$ with potentially unequal sizes, and it reduces to the conventional expression for $H(g)$ when subleading terms are neglected.
-It is an important part of our argument here, however, that the subleading terms be retained and carried through to the calculation of the mutual information.
-For the calculation of $H(g|c)$, the second term in Eq.
-(2), we use a similar three-part encoding, but one that takes advantage of the receiver’s knowledge of $c$ when communicating $g$.
-We first communicate $q_{g}$ as before, at the same information cost of $H(q_{g})=\log n$.
-In place of the vector $n^{(g)}$ we then communicate a contingency table $n^{(gc)}$, a matrix with elements $n_{rs}^{(gc)}$ equal to the number of objects that simultaneously belong to group $r$ in the candidate labeling $c$ and group $s$ in the ground truth $g$.
-Figure 1 shows an example of a contingency table for two labelings of the nodes in a small network.
-Figure 1: An example of a contingency table for two colorings of the same network, one with three colors and one with two.
-The entries in the $3\times 2$ contingency table $n^{(gc)}$ count the number of nodes that have each combination of colors.
-The row and column sums $n^{(g)}$ and $n^{(c)}$ of the contingency table count the number of nodes with each color in the two colorings.
-Note that, although we illustrate the contingency table with an application to a network, the table itself is independent of the network structure.
-Our process for transmitting the contingency table is analogous to the one we use for transmitting $n^{(g)}$.
-Since the labeling $c$ is known, the column sums $n^{(c)}$ of the contingency table are known as well, and we assume each column to be drawn from a Dirichlet-multinomial distribution with the appropriate sum.
-We use the same concentration parameter $\alpha_{g|c}$ for every column, but the columns are otherwise independent.
-If we denote column $s$ by $n_{\cdot s}^{(gc)}$, then the information cost to transmit the entire contingency table is
-$H($ $n^{(gc)}|n^{(c)},q_{g},\alpha_{g|c})=\sum_{s=1}^{q_{c}}H(n_{\cdot s}^{(gc)}|n_{s}^{(c)},q_{g},\alpha_{g|c})$
-$=\sum_{s=1}^{q_{c}}\biggl{[}\log\binom{n_{s}^{(c)}+q_{g}\alpha_{g|c}-1}{q_{g}\alpha_{g|c}-1}$
-$\hskip 60.00009pt-\sum_{r=1}^{q_{g}}\log\binom{n_{rs}^{(gc)}+\alpha_{g|c}-1}{\alpha_{g|c}-1}\biggr{]}.$ (9)
-This formulation allows for very efficient transmission of the contingency table under commonly occurring conditions.
-Consider for instance the special (but not implausible) case where $g=c$, so that $n^{(gc)}$ is diagonal.
-In this case the optimal concentration parameter is $\alpha_{g|c}=0$, which restricts the possible contingency tables to those for which each column $n_{\cdot s}^{(gc)}$ has exactly one nonzero entry.
-Given that we know the column sums, this leaves only the positions of the nonzero entries within each column to be transmitted, which results in near-optimal performance, a decisive improvement over previous reductions of the mutual information [7].
-Most practical applications will not have $c=g$ precisely, but the two labelings are often similar and similarly impressive improvements in performance can be realized in such cases as well.
-Finally, having transmitted the contingency table, it remains only to transmit the ground-truth labeling itself, where we need consider only those labelings consistent with the contingency table, given the known candidate labeling $c$.
-The number of such labelings is $\prod_{r}n_{r}^{(c)}!/\prod_{rs}n_{rs}^{(gc)}!$, so the information needed to uniquely identify one of them is
-$H(g|n^{(gc)},c)=\log\frac{\prod_{r}n_{r}^{(c)}!}{\prod_{rs}n_{rs}^{(gc)}!}.$ (10)
-In typical applications the number of labelings compatible with the contingency table is much smaller than the total number of labelings, and hence this encoding substantially reduces the amount of information needed to transmit the ground truth.
-It is not the only encoding possible, but it is an efficient one in practice and it is the one upon which the conventional mutual information is based.
-Collecting all our terms, the total conditional information can now be written
-$H(g|c)$ $=H(q_{g})+H(n^{(gc)}|n^{(c)},q_{g},\alpha_{g|c})+H(g|n^{(cg)},c).$ (11)
-And substituting (8) and (11) into Eq. (2), we get a complete expression for the mutual information
+I(c;g)=H(g)H(g|c).I(c;g)=H(g)-H(g|c). (1)
+Loosely, we say that $I(c;g)$ measures how much $c$ tells us about $g$. Since the conditional information content $H(g|c)$ is nonnegative, the mutual information must satisfy inequalities
+$I(c;g)\leq H(g),H(c),$ (2)
+which are saturated by
+$I(g;g)=H(g),\qquad I(c;c)=H(c).$ (3)
+Carefully accounting for all of the constituent information costs, we can compute the mutual information from [7]
 $I(c;g)$ $=H(g)-H(g|c)$
-$=\log\frac{n!\prod_{rs}n_{rs}^{(gc)}!}{\prod_{r}n_{r}^{(c)}!\prod_{s}n_{s}^{(g)}!}+\log\binom{n+q_{g}\alpha_{g}-1}{q_{g}\alpha_{g}-1}$
-$\qquad-\sum_{r=1}^{q_{g}}\log\binom{n_{r}^{(g)}+\alpha_{g}-1}{\alpha_{g}-1}$
-$\qquad-\sum_{s=1}^{q_{c}}\log\binom{n_{s}^{(c)}+q_{g}\alpha_{g|c}-1}{q_{g}\alpha_{g|c}-1}$
-$\qquad+\sum_{r=1}^{q_{g}}\sum_{s=1}^{q_{c}}\log\binom{n_{rs}^{(gc)}+\alpha_{g|c}-1}{\alpha_{g|c}+1}.$ (12)
-In most treatments, only the first of these terms is retained, since the others are subdominant and often small [6, 7].
-When all the terms are retained, however, the quantity (12) is known as the reduced mutual information and there are good reasons to use this full expression in certain circumstances, a point that we investigate in Section II.1.
-If we follow convention, however, and keep only the leading term, then we arrive at the traditional mutual information, which we denote $I_{0}$:
+$=\log\frac{n!\prod_{rs}n_{rs}^{(gc)}!}{\prod_{r}n_{r}^{(c)}!\prod_{s}n_{s}^{(g)}!}+\log\binom{n+q_{g}\alpha_{g}-1}{q_{g}\alpha_{g}-1}-\sum_{r=1}^{q_{g}}\log\binom{n_{r}^{(g)}+\alpha_{g}-1}{\alpha_{g}-1}$
+$\qquad-\sum_{s=1}^{q_{c}}\log\binom{n_{s}^{(c)}+q_{g}\alpha_{g|c}-1}{q_{g}\alpha_{g|c}-1}+\sum_{r=1}^{q_{g}}\sum_{s=1}^{q_{c}}\log\binom{n_{rs}^{(gc)}+\alpha_{g|c}-1}{\alpha_{g|c}+1}$. (4)
+In this expression the quantity $n_{r}^{(g)}$ denotes the number of objects in labeling $g$ that belong to group $r=1\ldots q_{g}$, and similarly for $n_{s}^{(c)}$.
+The quantities $n_{rs}^{(gc)}$ are the entries of the contingency table, and are equal to the number of objects that simultaneously belong to group $s$ in the candidate labeling $c$ and group $r$ in the ground truth $g$.
+Evaluating the sums and products over these table entries numerically, the value of $I(c;g)$ can be computed in time $\mathrm{O}(q_{c}q_{g})$.
+The parameter values $\alpha_{g}$ and $\alpha_{g|c}$ are chosen to optimize the transmission process using a fast line search, providing good compression for realistic labelings as discussed in [7].
+The first term in Eq.
+(4) represents the information cost of transmitting the labelings given knowledge of the contingency table $n_{rs}^{(gc)}$.
+The remaining terms account for the information costs associated with the transmission of the table itself.
+In some cases the information required to transmit the $q_{g}\times q_{c}$ contingency table is significantly smaller than that needed for the $n$-vectors representing the labelings, and it is neglected in most treatments [6, 7], leading to the traditional mutual information measure, which we denote $I_{0}$:
 $I_{0}(c;g)=\log\frac{n!\prod_{rs}n_{rs}^{(gc)}!}{\prod_{r}n_{r}^{(c)}!\prod_{s}n_{s}^{(g)}!}.$ (5)
-For later convenience, we also define entropies $H_{0}$ for the individual labelings ignoring subdominant terms, which can be thought of as the mutual information between the labelings and themselves:
-$H_{0}(c)$ $=I_{0}(c;c)=\log\frac{n!}{\prod_{r}n_{r}^{(c)}!},$ (14)
-$H_{0}(g)$ $=I_{0}(g;g)=\log\frac{n!}{\prod_{s}n_{s}^{(g)}!}.$ (15)
-Note that these differ from the entropy of Eq. (8), which includes subdominant terms. Trivially (14) and (15) provide upper bounds on the mutual information, while zero is a lower bound, so
-$0\leq I_{0}(c;g)\leq H_{0}(c),H_{0}(g).$ (16)
-These bounds will be useful shortly.
-A common further step is to approximate the mutual information of Eq. (13) using Stirling’s approximation in the form
-$\log n!=n\log n-n+\mathrm{O}(\log(n)),$ (17)
-which gives
-$I_{0}(c;g)=n\sum_{rs}p_{rs}^{(gc)}\log\frac{p_{rs}^{(gc)}}{p_{r}^{(c)}p_{s}^{(g)}}+\mathrm{O}\left(\log n\right),$ (18)
-where
-$p_{rs}^{(gc)}={n_{rs}^{(gc)}\over n},\qquad p_{r}^{(c)}={n_{r}^{(c)}\over n},\qquad p_{s}^{(g)}={n_{s}^{(g)}\over n},$ (19)
-and we adopt the convention that $00\log 0=$ 0.
-Equation (18) is often used as the definition of the mutual information, although technically it is only an approximation, albeit one that is asymptotically correct in the limit of large $n$.
-(It can also be derived by considering the information cost to transmit the true label of a single object, selected at random, with and without knowledge of the corresponding candidate label.)
-II.1 Reduced mutual information
-Equation (13) defines the standard mutual information.
-In this paper we explore variant definitions that can perform better at the clustering tasks we consider.
-A number of variants have been proposed, of which the simplest in the context of the current presentation is the reduced mutual information $I(c;g)$ of Eq.
-(12), which is simply the value of the mutual information when one does not neglect the subleading information cost of transmitting the group sizes and the contingency table.
-To understand the importance of the subleading terms let us explore how things can go wrong if we neglect them.
-To take an extreme example, suppose an algorithm simply places every object in a group of its own, which corresponds to the candidate labeling $c=(1,\ldots,n)$.
+In other cases, however, the information cost of the contingency table can play a significant role and should be retained as in Eq.
+(4), which yields the quantity known as the reduced mutual information.
+As an example, suppose a community detection algorithm simply places each node in its own group, resulting in a candidate labeling $c=(1,2,3,\ldots)$.
 No matter what the ground truth is, this choice of $c$ clearly contains no information about it whatsoever, so we might expect the mutual information to be zero.
-This, however, is not what we find.
-The contingency table in this case is
-$n_{rs}^{(gc)}=\biggl{\{}\begin{array}[]{ll}1&\quad\mbox{if $g_{r}=s$,}\\ 0&\quad\mbox{otherwise,}\\ \end{array}$ (20)
-so the conventional mutual information of Eq. (13) is
-$I_{0}(c;g)=\log\frac{n!}{\prod_{s}n_{s}^{(g)}!}=H_{0}(g).$ (21)
-This answer is not merely wrong; it is maximally so.
-The mutual information should take its minimum value of zero, but instead it takes the value $H_{0}(g)$, which is the maximum possible since $H_{0}(g)$ is an upper bound as we have said.
-The reason for this result is that in this case the contingency table itself uniquely defines $g$, so neglecting it puts the mutual information in error by an amount equal to the complete information cost of the ground truth.
-If we include the subleading terms on the other hand, this erroneous behavior disappears.
-Assuming for simplicity that the ground-truth groups are of equal size, the optimal concentration parameters are $\alpha_{g|c}=0$ and $\alpha_{g}=\infty$, and the reduced mutual information of Eq.
-(12) becomes
-$I(c;g)$ $=I_{0}(c;g)-\log\left[\frac{n!}{\prod_{r}n_{r}!}(1/q_{g})^{n}\right]-n\log q_{g}$
-$=0.$ (22)
-which is the correct answer.
-A related shortcoming of the standard mutual information is that for finite $n$ a random labeling $c$ will have positive mutual information with respect to any ground truth in expectation: because mutual information is non-negative, fluctuations due to randomness in finite systems will produce non-negative values only and hence their average will in general be positive [4, 9].
-This seems counterintuitive: we would expect the average value for a random labeling to be zero.
-We can rectify this issue by using another variant of the mutual information, which subtracts off the expected value of the measure, thereby making the average zero by definition.
-To do this we must first decide how a random labeling is to be defined—over what distribution of candidate labelings are we averaging?
+If we use the traditional mutual information, however, we get $I_{0}(c;g)=I_{0}(g;g)$, which is not merely nonzero, but implies that the candidate labeling is maximally informative, telling us everything about the ground truth.
+This is as wrong as it could be—in fact the candidate tells us nothing at all.
+The reduced mutual information, on the other hand, correctly says that $I(c;g)=0$ (see Section B in the Supplementary Materials).
+For the same reason, the traditional mutual information gives a maximum score to any candidate $c$ which is a refinement of the true $g$, i.e., every community in $c$ is a subset of a community in $g$.
+More generally, the traditional mutual information is biased towards labelings with too many groups [3, 4, 6, 7, 9].
+A related shortcoming of the traditional mutual information is that for finite $n$ even a random labeling $c$ will have positive mutual information with respect to any ground truth in expectation: because the traditional mutual information is non-negative, fluctuations due to randomness will produce non-negative values only and hence their average will in general be positive [4, 10].
+This seems counterintuitive; we would expect the average value for a random labeling to be zero.
+We can solve this problem by using another variant of the mutual information, which subtracts off the expected value, thereby making the average zero by definition.
+To do this we must first specify how the expectation is defined—over what distribution of candidate labelings are we averaging?
 The conventional choice is to take the uniform distribution over labelings that share the same group sizes $n^{(c)}$ as the actual candidate $c$.
@@
 [4]:
-$I_{A}(c;g)=I_{0}(c;g)-\langle I_{0}(c;g)\rangle_{\{c|n^{(c)}\}},$ (23)
-where the expectation is over the relevant ensemble.
-This measure can also be treated within the information-cost framework of the present paper, as described in [6].
+$I_{A}(c;g)=I_{0}(c;g)-\bigl\langle I_{0}(c;g)\bigr\rangle_{\{c|n^{(c)}\}}\,$, (6)
+where the expectation $\langle\ldots\rangle$ is over the relevant ensemble.
+The adjusted mutual information can also be derived in a fully information-theoretic manner, as described in [6].
 There it is shown that the subtracted term $\langle I_{0}(c;g)\rangle_{\{c|n^{(c)}\}}$ is precisely equal to the average cost of transmitting the contingency table when labelings are drawn from the uniform distribution.
-Thus, one can view the adjusted mutual information as a reduced measure, but differing from the one proposed in this paper in two ways.
-First, the correction term counts the cost of transmitting the contingency table but not the cost of transmitting the counts $n^{(g)}$, as described in Section II.
-This, however, is only a minor concern—the latter is usually much smaller than the former and its neglect introduces only a small error in most cases.
-The second difference is more serious and revolves around the choice to draw labelings uniformly from those with given $n^{(c)}$.
-As discussed in [6], this distribution heavily favors contingency tables with relatively uniform entries, simply because there are many more labelings that correspond to uniform tables than to non-uniform ones.
-Real contingency tables, on the other hand, are often highly non-uniform, since many applications of the mutual information focus on labelings that are somewhat similar to the ground truth (producing a non-uniform table).
+However, this distribution heavily favors contingency tables with relatively uniform entries, simply because there are many more labelings that correspond to uniform tables than to non-uniform ones.
+Real contingency tables, on the other hand, are often highly non-uniform, since applications of the mutual information focus on labelings that are somewhat similar to the ground truth (producing a non-uniform table).
 In such cases, the average used in the adjusted mutual information puts most of its weight on configurations that are very different from those that occur in reality, making it a poor representation of true information costs.
-The measure proposed in this paper, by contrast, deliberately allows for non-uniform tables by drawing them from a Dirichlet-multinomial distribution and we argue that this is a strong reason to favor it over the adjusted mutual information.
-In addition it is also derived naturally from a fully information-theoretic argument, by contrast with the more ad hoc derivation of the adjusted mutual information.
-Notwithstanding, in Section III we give results using both reduced and adjusted mutual information, and find fairly similar outcomes in the two cases.
-II.2 Normalization of the mutual information
+The reduced mutual information considered in this paper, by contrast, deliberately allows for non-uniform tables by drawing them from a Dirichlet-multinomial distribution and we argue that this is a strong reason to favor it over the adjusted mutual information.
+Nonetheless, in Section 2.3 we give results using both reduced and adjusted mutual information, and find fairly similar outcomes in the two cases.
+2.1 Normalization of the mutual information
 A fundamental difficulty with mutual information as a measure of similarity is that its range of values depends on the particular application, which makes it difficult to say when a value is large or small.
@@
 Sometimes it is and sometimes it isn’t, depending on the context.
-To get around this problem one commonly normalizes the mutual information so that it takes a maximum value of 1 when the candidate labeling agrees exactly with the ground truth.
-There are a number of ways this can be achieved but, as we show here, they are not all equal.
+To get around this obstacle one commonly normalizes the mutual information so that it takes a maximum value of 1 when the candidate labeling agrees exactly with the ground truth.
+There are a number of ways this can be achieved and, as we show here, they are not all equal.
 In particular, some, including the most popularly used normalization, can result in biased results and should, in our opinion, be avoided.
 In its place, we propose an alternative, unbiased normalized measure.
-The most popular normalized measure, commonly referred to simply as the “normalized mutual information,” uses the plain mutual information $I_{0}(c;g)$ as a base measure and normalizes it thus:
-$\text{NMI}_{0}^{(S)}(c;g)$ $=\frac{I_{0}(c;g)}{\frac{1}{2}[H_{0}(c)+H_{0}(g)]}$
-$=\frac{I_{0}(c;g)+I_{0}(g;c)}{I_{0}(c;c)+I_{0}(g;g)}.$ (24)
+The most popular normalized measure, commonly referred to simply as the normalized mutual information, uses the plain mutual information $I_{0}(c;g)$ as a base measure and normalizes it thus:
+$\text{NMI}_{0}^{(S)}(c;g)$ $=\frac{I_{0}(c;g)}{\frac{1}{2}[H_{0}(c)+H_{0}(g)]}=\frac{I_{0}(c;g)+I_{0}(g;c)}{I_{0}(c;c)+I_{0}(g;g)}.$ (7)
 This measure has a number of desirable features.
 Because of the inequalities in (2), its value falls strictly between zero and one.
-And since both the base measure and the normalization are symmetric under interchange of $c$ and $g$, the normalized measure also retains this symmetry (hence the superscript “$(S)$,” for symmetric).
-Figure 2: An example of how normalization can impact which labeling is preferred by a mutual information measure.
-For the ground-truth labeling $g$ (top), the standard unnormalized mutual information $I_{0}(c,g)$ of Eq.
-(18) gives different scores to the two candidate labelings $c_{A}$ and $c_{B}$, the former receiving a higher score than the latter.
-By definition, the asymmetrically normalized mutual information will always agree with the unnormalized measure, but the symmetrically normalized measure may not, as is the case here: in this example the symmetric measure scores candidate $c_{B}$ higher.
-Note that the network itself plays no role here—it is included only as a visual aid.
+And since both the base measure and the normalization are symmetric under interchange of $c$ and $g$, the normalized measure also retains this symmetry (hence the superscript $(S)$, for symmetric).
+Figure 1: Normalization can impact which labeling is preferred by a mutual information measure.
+In this example 64 objects are split into four equally sized true groups $g$, denoted by their color.
+Against this ground truth, we compare two candidate labelings, $c_{1}$ and $c_{2}$.
+The standard unnormalized mutual information $I_{0}(c;g)$ of Eq.
+(5) reports that labeling $c_{1}$ shares more bits of information with the ground truth than does $c_{2}$.
+By definition, the asymmetrically normalized mutual information $\text{NMI}^{(A)}$ will always agree with the unnormalized measure, as in this example.
+The symmetrically normalized $\text{NMI}^{(S)}$, on the other hand, is biased in favor of simpler labelings, which makes it prefer labeling $c_{2}$ in this case.
 Equation (7) is not the only normalization that achieves these goals. Equation (2) implies that
-$I_{0}(c;g)$ $\leq\min(I_{0}(c;c),I_{0}(g;g))$ (25)
-$\leq\sqrt{I_{0}(c;c)I_{0}(g;g)}$ (26)
+$I_{0}(c;g)$ $\leq\min(I_{0}(c;c),I_{0}(g;g))$
+$\leq\sqrt{I_{0}(c;c)I_{0}(g;g)}$
 $\leq\max(I_{0}(c;c),I_{0}(g;g)),$ (8)
 which gives us three more options for a symmetric denominator in the normalized measure. The arithmetic mean in Eq. (7), however, sees the most use by far [11, 12, 13, 14].
-We can extend the notion of symmetric normalization to any other base measure of mutual information $I_{X}(c;g)$ by writing
+We can extend the notion of symmetric normalization to any other base measure of mutual information $I_{X}(c;g)$ satisfying the inequality Eq.
+(2), such as adjusted or reduced mutual information [7], by writing
 $\text{NMI}_{X}^{(S)}(c;g)=\frac{I_{X}(c;g)+I_{X}(g;c)}{I_{X}(c;c)+I_{X}(g;g)}.$ (9)
 All such measures, however, including the standard measure of Eq.
-(24), share a crucial shortcoming, namely that the normalization depends on the candidate labeling $c$ and hence that the normalized measure can prefer a different candidate labeling to the base measure purely because of the normalization.
+(7), share a crucial shortcoming, that the normalization depends on the candidate labeling $c$ and hence that the normalized measure can prefer a different candidate labeling to the base measure purely because of the normalization.
 Figure 1 shows an example of how this can occur.
-Two candidate labelings $c_{A}$ and $c_{B}$ are considered for the same ground truth $g$.
+In this example 64 objects are split into four equally sized groups in the ground-truth labeling $g$, against which we compare two candidate labelings, $c_{1}$ and $c_{2}$.
+Neither candidate is perfect.
+Labeling $c_{1}$ identifies the correct number of groups but misplaces certain objects, while labeling $c_{2}$ contains only two groups, each an aggregation of a pair of true groups.
 Under the unnormalized mutual information of Eq.
-(13), candidate A receives a higher score than candidate B, but under the normalized measure of Eq.
+(5), $c_{1}$ receives a higher score than $c_{2}$, but under the normalized measure of Eq.
 (7) the reverse is true.
-This behavior is due to the difference in entropy between the two candidate divisions, the one on the right having larger entropy than the one on the left, which increases its normalization factor and correspondingly decreases the normalized mutual information.
+This behavior is due to the difference in entropy $I_{0}(c;c)$ between the two candidate divisions.
+Since the candidate entropy appears in the denominator of the symmetric normalization, Eq.
+(7), the symmetric measure favors simple labelings with smaller entropy.
+In this example, candidate $c_{2}$ has only two groups, which results in a smaller entropy and a bias in its favor.
 We argue that the unnormalized measure is more correct on this question, having a direct justification in terms of information theory.
-The purpose of the normalization is purely to map the values of the measure onto a convenient numerical interval, and should not change outcomes as it does here.
+The purpose of the normalization is merely to map the values of the measure onto a convenient numerical interval, and should not change the outcome as it does here.
 Moreover, different symmetric normalizations can produce different results.
 For instance, if one normalizes by $\max(I_{0}(c;c),I_{0}(g;g))$ in Fig.
-2 then candidate $c_{A}$ is favored in all cases.
+1 then candidate $c_{1}$ is favored in all cases.
+The choice of normalization can also change conclusions even when the number of groups does not vary.
+Consider the example shown in Fig.
+2, in which the ground truth $g$ splits the objects into two equally sized groups.
+Candidate $c_{1}$ also has two groups of equal size, although each is slightly polluted with the objects of the opposite group.
+Candidate $c_{2}$ has two groups of different sizes, a small group with one color only and a larger group with objects of both colors.
+Although both labelings have the same number of groups, candidate $c_{2}$ has a smaller entropy $I_{0}(c;c)$, since it is easier to communicate group identity within $c_{2}$—most objects belong to the larger group.
+As a result, the symmetrically normalized mutual information prefers $c_{2}$, while the unnormalized measure prefers $c_{1}$.
+Figure 2: Another example where the choice of normalization affects which labeling is preferred, this one with the same number of groups in both divisions.
+Although both candidates $c_{1}$ and $c_{2}$ correctly split the objects into two groups, candidate labeling $c_{1}$ scores higher with respect to both the unnormalized and asymmetrically normalized mutual informations while the symmetric normalization prefers $c_{2}$.
 These issues are unavoidable when using a symmetric normalization scheme.
 In any such scheme the normalization must depend on both $c$ and $g$ and hence can vary with the candidate labeling.
-However, if we drop the requirement of symmetry then we can normalize in a straightforward way that avoids these issues.
+However, if we drop the requirement of symmetry then we can normalize in a way that avoids these issues.
 We define the asymmetric normalization of any base measure $I_{X}$ as
 $\text{NMI}_{X}^{(A)}(c;g)=\frac{I_{X}(c;g)}{I_{X}(g;g)}.$ (10)
-This definition still gives $I_{X}^{(A)}(g;g)=1$, but now the normalization factor in the denominator has no effect on choices between candidate labelings, since it is independent of $c$.
+This definition still gives $\text{NMI}_{X}^{(A)}(g;g)=1$, but now the normalization factor in the denominator has no effect on choices between candidate labelings, since it is independent of $c$.
 In fact, Eq.
@@
 Thus this measure also removes any ambiguity about how one should perform the normalization.
-Loosely, this asymmetrically normalized mutual information measures “how much $c$ tells us about $g$ as a fraction of all there is to know about $g$.”
-The amount of bias inherent in the symmetrically normalized measure when compared with asymmetric one can be quantified by the ratio between the two:
+Loosely, this asymmetrically normalized mutual information measures how much $c$ tells us about $g$ as a fraction of all there is to know about $g$.
+Asymmetrically normalized measures have appeared in other contexts under the name “coefficient of constraint” or “coefficient of uncertainty” [15, 16, 17], although symmetric normalizations are almost universally used in the model validation context, where they introduce the bias we discuss here.
+The amount of bias inherent in the symmetrically normalized measure when compared with the asymmetric one can be quantified by the ratio between the two:
 ${\text{NMI}_{X}^{(S)}(c;g)\over\text{NMI}_{X}^{(A)}(c;g)}$ $={[I_{X}(c;g)+I_{X}(g;c)]/[I_{X}(c;c)+I_{X}(g;g)]\over I_{X}(c;g)/I_{X}(g;g)}$
@@
 ${\text{NMI}_{X}^{(S)}(c;g)\over\text{NMI}_{X}^{(A)}(c;g)}={H(g)\over\frac{1}{2}[H(g)+H(c)]}.$ (12)
-Note how this expression results in a particular bias against complex candidate labelings $c$ that have higher entropy $H(c)$ than the ground truth.
-Traditional mutual information measures are particularly problematic when the candidate is a refinement of the ground truth, meaning the candidate groups are subsets of the ground-truth groups.
-The example given in Section II.1, in which the candidate labeling puts every node in a group on its own, is an extreme instance of this.
-Equation (31) shows that, to some extent, the traditional symmetric normalization corrects this issue: a candidate $c$ that takes the form of a refinement of $g$ will have $H(c)>H(g)$ and hence the normalized mutual information will be penalized.
+Values of this quantity below (above) 1 indicate that the symmetric measure is biased low (high).
+Thus, for instance, complex candidate labelings $c$ that have higher entropy than the ground truth will result in a symmetric measure whose values are too low.
+As discussed in the introduction, traditional mutual information measures are particularly problematic when the candidate is a refinement of the ground truth, meaning the candidate groups are subsets of the ground-truth groups.
+In that case the traditional measure returns values that are too high.
+Equation (12) implies that, to some extent, the symmetric normalization will correct this issue: a candidate $c$ that takes the form of a refinement of $g$ will have $H(c)>H(g)$, which will lower the value of the symmetrically normalized mutual information.
+This may in part explain why these biases have been overlooked in the past: two wrongs have conveniently canceled to (approximately) make a right.
 We argue, however, that this is not the best way to address the problem and that the correct approach is instead to use the reduced mutual information.
-The reduced mutual information also penalizes such cases, but it does so in a more principled manner that directly addresses the root cause of the problem, rather than merely penalizing complex candidate labelings in an ad hoc manner as a side-effect of normalization.
-We will see examples of this in Section III, where we apply both measures to community detection in networks and find that, although the symmetrically normalized traditional measure is biased, it is in some cases fortuitously biased in the right direction, although it is still problematic in some others.
+The reduced mutual information also corrects for the case where one labeling is a refinement of the other, but does so in a more principled manner that directly addresses the root cause of the problem, rather than merely penalizing complex candidate labelings in an ad hoc manner as a side-effect of normalization.
+We will see examples of these effects in Section 2.2, where we apply the various measures to community detection in networks and find that indeed the traditional symmetrically normalized mutual information is biased.
+In some cases it is fortuitously biased in the right direction, although it is still problematic in some others.
 An obvious downside of asymmetric normalization is the loss of the symmetry in the final measure.
-In the most common applications of normalized mutual information, where labelings are evaluated against a ground truth, an inherently asymmetric situation, the asymmetric measure makes perfect sense, but in other cases the lack of symmetry can be undesirable.
+In the most common applications of normalized mutual information, where labelings are evaluated against a ground truth, an inherently asymmetric situation, the asymmetric measure makes sense, but in other cases the lack of symmetry can be undesirable.
 Embedding and visualization methods that employ mutual information as a similarity measure, for example, normally demand symmetry [18].
-And in cases where one is comparing two candidate labelings directly to one another, rather than to a separate ground truth, then a symmetric measure may be preferable.
-Even in this case, however, the asymmetric measure may sometimes be the better choice, as discussed in the introduction.
+And in cases where one is comparing two candidate labelings directly to one another, rather than to a separate ground truth, a symmetric measure may be preferable.
+Even in this latter case, however, the asymmetric measure may sometimes be the better choice, as discussed in the introduction.
 For instance, when one labeling $c_{1}$ is a refinement of the other $c_{2}$, the information content is inherently asymmetric: $c_{1}$ says more about $c_{2}$ than $c_{2}$ does about $c_{1}$.
@@
 Because of this nested relationship, it is relatively easy to transmit $c_{2}$ given knowledge of $c_{1}$ but more difficult to do the reverse.
-This imbalance is reflected in the asymmetrically normalized mutual information values in each direction (top and bottom arrows in the figure), but absent from the symmetric version (middle row).
+This imbalance is reflected in the asymmetric normalized mutual information values in each direction (top and bottom arrows in the figure), but absent from the symmetric version (middle row).
 Combining the benefits of asymmetric normalization and the reduced mutual information, we advocate in favor of the asymmetrically-normalized reduced mutual information defined by
@@
 (4).
-This measure correctly accounts for the information contained in the contingency table, returns 0 in expectation when $c$ is uncorrelated with the ground truth and 1 when $c=g$, and always favors the same labeling as the unnormalized measure.
-Figure 3: Values of the various normalized mutual information measures we consider, between two example partitions of the same set of 27 objects.
-The left partition $c_{1}$ divides the objects into nine groups of three objects each and is a refinement of the right partition $c_{2}$, which divides them into only three groups, aggregating those on the left.
+This measure correctly accounts for the information contained in the contingency table, returns a negative value when $c$ is unhelpful for recovering the ground truth, returns 1 if and only if $c=g$, and always favors the same labeling as the unnormalized measure.
+The traditional normalized mutual information possesses none of these desirable qualities.
+Figure 3: Values of the various normalized mutual information measures between two example partitions of the same set of 27 objects.
+The left partition $c_{1}$ divides the objects into nine groups of three objects each and is a refinement of the right partition $c_{2}$, which divides them into only three groups.
 The arrows indicate the direction of the comparison and the accompanying notations give the values of the mutual information measures.
 All three asymmetrically normalized measures capture the intuition that the partition $c_{1}$ tells us more about $c_{2}$ than $c_{2}$ tells us about $c_{1}$.
-III Example application:
-Community detection
-As an example of the performance of the various measures discussed above, we describe in this section the results of an extensive series of tests in which the measures are used to score the output of algorithms for network community detection.
-In these tests we use the popular Lancichinetti-Fortunato-Radicchi (LFR) graph model [15] to generate random test networks with known community structure and realistic distributions of node degrees and group sizes, and then attempt to recover that structure using a variety of standard community detection algorithms, quantifying the accuracy of the recovery with six different measures: symmetrically normalized versions of the traditional mutual information, the adjusted mutual information, and the reduced mutual information, and asymmetrically normalized versions of the same.
+2.2 Copurchasing network of books
+As an example of the performance of the various measures discussed above, in this section we use them to score the output of algorithms for network community detection.
+We consider a selection of networks with known true groups and attempt to recover those groups using various standard community detection algorithms, quantifying the accuracy of recovery with six different measures: the symmetrically and asymmetrically normalized versions of the traditional mutual information, the adjusted mutual information, and the reduced mutual information.
+As a first example, Fig.
+4 shows a network compiled by V.
+Krebs of 105 books about US politics published around the time of the 2004 US presidential election.
+The true groups $g$ in this case represent the political lean of each book’s content: liberal, conservative, or neutral [19], as determined by Krebs.
+Network edges connect pairs of books frequently purchased together on Amazon.com, and such books typically are of the same political bent, either both liberal or both conservative, so that the political positions form communities of connected network nodes.
+Community detection algorithms, which aim to find such communities, should therefore be able to recover political positions, at least approximately, from the network of connections alone.
+The figure shows the results of applying two of the most popular such algorithms, InfoMap and modularity maximization (with resolution parameter $\gamma=2$), but we find that these algorithms identify quite different partitions of the network, with InfoMap finding six groups where modularity finds nine.
+We use the normalized mutual information in its various forms to assess how closely each division matches the ground truth.
+We find that both the base mutual information measure $I_{X}$ and its normalization considerably impact our assessments.
+For instance, the asymmetrically normalized traditional mutual information gives a score of $\text{NMI}_{0}^{(A)}(c_{2};g)=0.728$, out of a maximum of 1, for the partition found by modularity maximization, which is over twice the score of $\text{NMI}^{(S)}(c_{2};g)=0.294$ given to the same partition by the symmetrically normalized reduced measure.
+The choice of mutual information measure can also change which algorithm is preferred.
+For the traditional unreduced base measure $I_{0}$, the asymmetrically normalized mutual information favors the groups found by modularity while the symmetric version favors the simpler partition found by InfoMap.
+The same pattern holds if the adjusted $I_{A}$ is used.
+The reduced measure $I$, however, prefers InfoMap regardless of normalization, although by a narrower margin in the asymmetric case.
+As discussed in the previous section, we favor the asymmetrically normalized reduced measure on formal grounds, and view the (often large) deviations from its assessment as unnecessary biases.
+Figure 4: Performance of the InfoMap and $\gamma=2$ modularity maximization algorithms in recovering the true groups in the copurchasing network of political books [19].
+True and recovered groups are indicated by node color, and similarity between the two is assessed using the mutual information in its traditional $I_{0}$, adjusted $I_{A}$, and reduced II variants, normalized asymmetrically $\text{NMI}^{(A)}$ and symmetrically $\text{NMI}^{(S)}$.
+The choices of both the base measure and the normalization impact which algorithm is preferred.
+2.3 LFR benchmark networks
+In one sense the previous example is unusual: there are relatively few cases like this of networks with known ground-truth communities to compare against.
+To get a more comprehensive picture of algorithm performance, therefore, researchers have turned to synthetic benchmark networks like those generated by the popular Lancichinetti-Fortunato-Radicchi (LFR) graph model [20], which creates networks with known community structure and realistic distributions of node degrees and group sizes.
 A number of studies have been performed in the past to test the efficacy of community detection algorithms on LFR benchmark networks [11, 12, 13, 14], but using only the symmetrically normalized, non-reduced mutual information as a similarity measure.
@@
 The LFR model contains a number of free parameters that control the size of the networks generated, their degree distribution, the distribution of community sizes, and the relative probability of within- and between-group edges.
-(We give further details on the LFR generative process in Appendix A.) We find that the distributions of degrees and community sizes do not significantly impact the relative performance of the various algorithms tested and that performance differences are driven primarily by the size $n$ of the networks and the mixing parameter $\mu$ that controls the ratio of within/between-group connections, so our tests focus on performance as a function of these variables.
-Figure 4: The colors in each panel in the center of this figure indicate which of three community detection algorithms (InfoMap (red) and modularity maximization with $\gamma=1$ (green) and $\gamma=10$ (blue)—see key on right) are best able to find the known communities in a large set of LFR benchmark networks, according to the six mutual information measures we consider.
+(See Section A in the Supplementary Materials for details of the LFR generative process.) We find that the distributions of degrees and community sizes do not significantly impact the relative performance of the various algorithms tested and that performance differences are driven primarily by the size $n$ of the networks and the mixing parameter $\mu$ that controls the ratio of connections within and between groups, so our tests focus on performance as a function of these parameters.
+Figure 5: Performance of three community detection algorithms: InfoMap (red) and modularity maximization with $\gamma=1$ (green) and $\gamma=10$ (blue)—see key on right.
+The colors in each panel of this figure indicate which of the three is best able to find the known communities in a large set of LFR benchmark networks, according to the six mutual information measures we consider.
 Mixtures of red, green, and blue denote the proportions of test cases in which each algorithm performs best.
 Regions in gray indicate parameter values for which no algorithm achieved a positive mutual information score.
-We perform community detection using six well-known algorithms as follows.
-(We use the implementations found in the igraph library [16], except for the inference method, for which we use the graph-tool library [17].)
+2.4 Comparison between variants of the mutual information
+Figure 5 summarizes the relative performance of the various mutual information measures in our tests.
+In this set of tests we limit ourselves, for the sake of clarity, to the top three community detection algorithms—InfoMap and the two variants of modularity maximization—and measure which of the three returns the best results according to each of our six mutual information measures, as a function of network size $n$ and the mixing parameter $\mu$.
+Each point in each of the six panels is color-coded with some mix of red, green, and blue to indicate in what fraction of cases each of the algorithms performs best according to each of the six measures and, as we can see, the results vary significantly among measures.
+An experimenter trying to choose the best algorithm would come to substantially different conclusions depending on which measure they use.
+One consistent feature of all six mutual information measures is the large red area in each panel of Fig.
+5, which represents the region in which the InfoMap algorithm performs best.
+Regardless of the measure used, InfoMap is the best performer on networks with low mixing parameter (i.e., strong community structure) and relatively large network size.
+For higher mixing (weaker structure) or smaller network sizes, modularity maximization does better.
+Which version of modularity is best, however, depends strongly on the mutual information measure.
+The traditional symmetrically normalized mutual information (top left panel) mostly favors the version with a high resolution parameter of $\gamma=10$ (blue), but the asymmetric reduced measure for which we advocate (bottom right) favors the version with $\gamma=1$ (green).
+(The regions colored gray in the figure are those in which no algorithm receives a positive mutual information score and hence all algorithms can be interpreted as failing.)
+These results raise significant doubts about the traditional measure.
+Consider the lower right corner of each plot in Fig.
+5, which is the regime of small network sizes $n$ and high mixing $\mu$ so that the community signal is weak and the noise is high.
+Here, the adjusted and reduced mutual informations indicate that all algorithms are failing.
+This is expected: for very weak community structure all detection algorithms are expected to show a “detectability threshold” beyond which they are unable to identify any communities [21, 22, 23].
+The standard normalized mutual information, on the other hand, claims to find community structure in this regime using the $\gamma=10$ version of modularity maximization.
+This occurs because the $\gamma=10$ algorithm finds many small communities and, as discussed in Section 2, a labeling with many communities, even completely random ones, is accorded a high score by a non-reduced mutual information.
+The presence of this effect is demonstrated explicitly in the Supplementary Materials, Section A.1 for the case of modularity maximization with $\gamma=10$, and this offers a clear reason to avoid the standard measure.
+Figure 6: Performance of each of the six community detection algorithms considered here in identifying the known communities in a large set of LFR benchmark networks, as quantified by the asymmetrically normalized reduced information measure $I^{(A)}(c;g)$ proposed in this paper.
+Gray areas indicate parameter values for which $\text{NMI}<0$, so that the conditional encoding is less efficient than the direct encoding.
+The bottom left panel in Fig.
+5 shows results for the asymmetrically normalized version of the traditional mutual information, which gives even worse results than the symmetric version, with hardly any region in which the $\gamma=1$ version of modularity maximization outperforms the $\gamma=10$ version.
+This behavior arises for the reasons discussed in Section 2.1—the bias inherent in the symmetric normalization fortuitously acts to partially correct the errors introduced by neglecting the information content of the contingency table.
+The asymmetric normalization eliminates this correction and hence performs more poorly.
+The correct solution to this problem, however, is not to use a symmetric normalization, which can bias outcomes in other ways as we have seen, but rather to adopt a reduced mutual information measure.
+Finally, comparing the middle and right-hand columns of Fig.
+5, we see that the results for the adjusted and reduced mutual information measures are quite similar in these tests, although there are some differences.
+In particular, the adjusted measure appears to find more significant structure for higher mixing than the reduced measure.
+This occurs because, as discussed in Section 2 and Ref.
+[6], the adjusted measure encodes the contingency table in a way that is optimized for more uniform tables than the reduced measure, and thus penalizes uniform tables less severely, leading to overestimates of the mutual information in the regime where detection fails—in this regime the candidate and ground-truth labelings are uncorrelated which results in a uniform contingency table.
+This provides further evidence in favor of using a reduced mutual information measure.
+2.5 Comparison between community detection algorithms
+Settling on the asymmetrically normalized reduced mutual information as our preferred measure of similarity, we now ask which community detection algorithm or algorithms perform best according to this measure?
+We have already given away the answer—InfoMap and modularity maximization get the nod—but here we give evidence for that conclusion.
+Figure 6 shows results for all six algorithms listed in the Supplementary Materials, Section A.2.
+Examining the figure, we see that in general the best-performing methods are InfoMap, traditional modularity maximization with $\gamma=1$, and the inference method using the degree-corrected stochastic block model.
+Among the algorithms considered, InfoMap achieves the highest mutual information scores for lower values of the mixing parameter $\mu$ in the LFR model, but fails abruptly as $\mu$ increases, so that beyond a fairly sharp cutoff around $\mu=0.5$ other algorithms do better.
+As noted by previous authors [24], InfoMap’s specific failure mode is that it places all nodes in a single community and this behavior can be used as a simple indicator of the failure regime.
+In this regime one must use another algorithm.
+Either the modularity or inference method are reasonable options, but modularity has a slight edge, except in a thin band of intermediate $\mu$ values which, in the interests of simplicity, we choose to ignore.
+(We discuss some caveats regarding the relationship between the degree-corrected stochastic block model and the LFR benchmark in Section A of the Supplementary Materials.)
+Thus—always assuming the LFR benchmark is a good test of performance—our recommendations for the best community detection algorithm are relatively straightforward.
+If we are in a regime where InfoMap succeeds, meaning it finds more than one community, then one should use InfoMap.
+If not, one should use standard modularity maximization with $\gamma=1$.
+That still leaves open the question of how the modularity should be maximized.
+In our studies we find the best results with simulated annealing, but simulated annealing is computationally expensive.
+In regimes where it is not feasible, we recommend using the Leiden algorithm instead.
+(Tests using other computationally efficient maximization schemes, such as the Louvain and spectral algorithms, generally performed less well than the Leiden algorithm.)
+3 Discussion
+In this paper we have examined the performance of a range of mutual information measures for comparing labelings of objects in classification, clustering, or community detection applications.
+We argue that the commonly used normalized mutual information is biased in two ways: (1) because it ignores the information content of the contingency table, which can be large, and (2) because the symmetric normalization it employs introduces spurious dependence on the labeling.
+We argue in favor of a different measure, an asymmetrically normalized version of the reduced mutual information, which rectifies both of these shortcomings.
+To demonstrate the effects of using different mutual information measures, we have presented results of an extensive set of numerical tests on popular network community detection algorithms, as evaluated by the various measures we consider.
+We find that conclusions about which algorithms are best depend substantially on which measure we use.
+4 Data availability
+All datasets analyzed here have been deposited in the Zenodo database under accession code 10.5281/zenodo.17211478, except for the synthetic LFR benchmark networks, which were randomly generated using the open-source NetworkX package as described in the Supplementary Materials, Section A.
+Code availability
+Python code implementing the mutual information measures discussed in this paper is available at https://github.com/maxjerdee/clustering-mi or [25].
+Acknowledgments
+The authors thank Samin Aref for useful comments and feedback.
+This work was supported in part by the US National Science Foundation under grants DMS–2005899 and DMS–2404617 (MN), the National Science Foundation of China through Young Scientist Fund Project grant 12405044 (AK), and computational resources provided by the Advanced Research Computing initiative at the University of Michigan.
+Author contributions
+All authors (MJ, AK, and MN) conceptualized the research and wrote and edited the paper. Code development and data analysis were performed by MJ. All authors have read the final manuscript.
+Competing interests
+The authors declare no competing interests.
+Appendix A Benchmarking
+In this appendix we give some additional details of our community detection tests.
+A.1 Numbers of groups
+As discussed in Section 2.4, the traditional (non-reduced) mutual information favors the $\gamma=10$ generalized modularity maximization algorithm over other algorithms across many of the tests reported in Fig.
+5.
+However, a closer inspection of the partitions found by this algorithm reveals that it drastically overestimates the number of communities.
+The number of groups inferred by each algorithm for each value of the mixing parameter $\mu$ is shown for $n=3200$ in Fig.
+7.
+The traditional mutual information has a bias towards labelings with an excessive number of groups, which causes it to prefer the $\gamma=10$ algorithm in this regime, while the reduced and adjusted mutual information measures prefer simple modularity maximization ($\gamma=1$).
+Figure 7: The number of groups inferred by each of the six algorithms in Fig.
+5 for LFR benchmark networks with $n=3200$ nodes and a range of values of the mixing parameter $\mu$.
+The ground-truth number of groups is shown in black.
+The InfoMap algorithm (red) generates an accurate number of groups for values of $\mu$ up to about 0.5, but beyond this point it erroneously places all nodes in a single group.
+Standard modularity maximization with resolution parameter $\gamma=1$ (green) underestimates the number of groups, perhaps because of the resolution limit on the detection of small groups, but not as severely as the number is overestimated when $\gamma=10$ (blue).
+A.2 Community detection algorithms
+We assess the performance of six common community detection algorithms on synthetic network examples.
+We use the algorithm implementations from the igraph library [26], except for the inference method, for which we use the graph-tool library [27].
+The six algorithms are as follows.
   1. 1.
@@
   2. 2.
-    Modularity maximization: Modularity is a quality function for network labelings equal to the fraction of edges within communities minus the expected such fraction if edge positions are randomized while preserving the node degrees.
+    Modularity maximization: Modularity is a quality function for network community divisions equal to the fraction of edges within communities minus the expected such fraction if edge positions are randomized while preserving node degrees.
 The labeling with the highest modularity is considered the best community division.
-Modularity can be maximized using a number of heuristic methods, of which the most popular are agglomerative methods such as the Louvain and Leiden algorithms [19, 20], spectral methods [21], and simulated annealing [22, 23, 24].
+Exact modularity maximization is NP-hard and usually intractable, but modularity can be maximized approximately using various heuristics, of which the most popular are agglomerative methods such as the Louvain and Leiden algorithms [29, 30], spectral methods [31], and simulated annealing [32, 33, 34].
 In our tests we use simulated annealing where computationally feasible and the Leiden algorithm otherwise, these approaches giving the most consistent maximization of the modularity.
@@
     Labelprop: The label propagation or “labelprop” algorithm initially places every node in its own community then iteratively updates the labels of randomly chosen nodes by majority vote among their network neighbors, breaking ties at random [39].
-As described in Section III.2, all of these algorithms perform reasonably well, but, as we will see, the best performers in the context of the current tests are InfoMap and the two variants of modularity maximization.
-III.1 Comparison between variants of the
-mutual information
-Figure 4 summarizes the relative performance of the various mutual information measures in our tests.
-In this set of tests we limited ourselves, for the sake of clarity, to the top three community detection algorithms—InfoMap and the two variants of modularity maximization—and measured which of the three returned the best results according to each of our six mutual information measures, as a function of network size $n$ and the mixing parameter $\mu$.
-Each point in each of the six panels in the figure is color-coded with some mix of red, green, and blue to indicate in what fraction of cases each of the algorithms performs best according to each of the six measures and, as we can see, the results vary significantly among measures.
-An experimenter trying to choose the best algorithm would come to substantially different conclusions depending on which measure they use.
-One consistent feature of all six mutual information measures is the large red area in each panel of Fig.
-4, which represents the region in which InfoMap performs best.
-Regardless of the measure used, InfoMap is the best performer on networks with low mixing parameter (i.e., strong community structure) and relatively large network size.
-For higher mixing (weaker structure) or smaller network sizes, modularity maximization does better.
-Which version of modularity is best, however, depends strongly on the mutual information measure.
-The traditional symmetrically normalized mutual information (top left panel) mostly favors the version with a high resolution parameter of $\gamma=10$ (blue), but the asymmetric reduced measure for which we advocate (bottom right) favors the version with $\gamma=1$ (green).
-(The regions colored gray in the figure are those in which no algorithm receives a positive mutual information score and hence all algorithms can be interpreted as failing.)
-These results raise significant doubts about the traditional measure.
-At large $n$ and high mixing $\mu$ the network will have many small communities, and in this regime it is natural that modularity maximization with a high resolution parameter will perform well because it favors large numbers of communities, and all mutual information measures agree on this point.
-For smaller system sizes or lower mixing, however, the network will have fewer, larger communities and in this regime we find that the $\gamma=10$ algorithm dramatically overestimates the number of communities in the network.
-With a non-reduced mutual information there is no penalty for doing this, as discussed in Section II.1, which is why the traditional normalized mutual information ranks the $\gamma=10$ algorithm highly in this regime, arguably erroneously.
-The reduced mutual information, by contrast, prefers the $\gamma=1$ algorithm.
-The number of groups inferred by each algorithm for each value of the mixing parameter $\mu$ is shown for $n=3200$ in Figure 5, which highlights these findings.
-Figure 5: The number of groups inferred by each of the six algorithms in Fig.
-4 for LFR benchmark networks with $n=3200$ nodes and a range of values of the mixing parameter $\mu$.
-The true number of groups is shown in black.
-The InfoMap algorithm (red) generates an accurate number of groups for values of $\mu$ up to about 0.5, but beyond this point it erroneously places all nodes in a single group.
-Standard modularity maximization with resolution parameter $\gamma=1$ (green) underestimates the number of groups, presumably because of the resolution limit on the detection of small groups, but not as severely as the number is overestimated when $\gamma=10$ (blue).
-An extreme example of the disagreement between measures can be seen in the lower right corner of each plot in Fig.
-4, where most measures return a negative value, indicating that community detection has failed altogether.
-This is expected: for very weak community structure all detection algorithms are expected to show a “detectability threshold” beyond which they are unable to identify any communities [30, 31, 32].
-The standard normalized mutual information, however, claims to find community structure in this regime using the $\gamma=10$ version of modularity maximization.
-This occurs because the $\gamma=10$ algorithm finds many small communities and, as discussed in Section II.1, a labeling with many communities, even completely random ones, is accorded a high score by a non-reduced mutual information.
-This offers a clear reason to avoid the standard measure.
-Figure 6: Performance of each of the six community detection algorithms considered here in identifying the known communities in a large set of LFR benchmark networks, as quantified by the asymmetrically normalized reduced information measure $I^{(A)}(c;g)$ proposed in this paper.
-The bottom left panel in Fig.
-4 shows results for the asymmetrically normalized version of the traditional mutual information, which gives even worse results than the symmetric version, with hardly any region in which the $\gamma=1$ version of modularity maximization outperforms the $\gamma=10$ version.
-This behavior arises for the reasons discussed in Section II.2: the bias inherent in the symmetric normalization fortuitously acts to partially correct the errors introduced by neglecting the information content of the contingency table.
-The asymmetric normalization eliminates this correction and hence performs more poorly.
-The correct solution to this problem, however, is not to use a symmetric normalization, which can bias outcomes in other ways as we have seen, but rather to adopt a reduced mutual information measure.
-Finally, comparing the middle and right-hand columns of Fig.
-4, we see that the results for the adjusted and reduced mutual information measures are quite similar in these tests, although there are some differences.
-In particular, the adjusted measure appears to find more significant structure for higher mixing than the reduced measure.
-This occurs because, as discussed in Section II.1 and Ref.
-[6], the adjusted measure encodes the contingency table in a way that is optimized for more uniform tables than the reduced measure, and thus penalizes uniform tables less severely, leading to overestimates of the mutual information in the regime where detection fails—in this regime the candidate and ground-truth labelings are uncorrelated which results in a uniform contingency table.
-This provides further evidence in favor of using a reduced mutual information measure.
-III.2 Comparison between
-community detection algorithms
-Settling on the asymmetrically normalized reduced mutual information as our preferred measure of similarity, we now ask which community detection algorithm or algorithms perform best according to this measure?
-We have already given away the answer—InfoMap and modularity maximization get the nod—but here we give evidence for that conclusion.
-Figure 6 shows results for all six algorithms listed in Section III.
-Examining the figure, we see that in general the best-performing methods are InfoMap, traditional modularity maximization with $\gamma=1$, and the inference method using the degree-corrected stochastic block model.
-Among the algorithms considered, InfoMap achieves the highest mutual information scores for lower values of the mixing parameter $\mu$ in the LFR model, but fails abruptly as $\mu$ increases, so that beyond a fairly sharp cutoff around $\mu=0.5$ other algorithms do better.
-As noted by previous authors [33], InfoMap’s specific failure mode is that it places all nodes in a single community and this behavior can be used as a simple indicator of the failure regime.
-In this regime one must use another algorithm.
-Either modularity maximization or inference using the degree-corrected stochastic block model are reasonable options, but modularity has a slight edge, except in a thin band of intermediate $\mu$ values which, in the interests of simplicity, we choose to ignore.
-(We discuss some caveats regarding the relationship between the degree-corrected stochastic block model and the LFR benchmark in Appendix A.)
-Thus our recommendations for the best community detection algorithm are relatively straightforward, if we base our choice on performance under the LFR benchmark, as is commonly done in this field.
-If we are in a regime where InfoMap succeeds, meaning it finds more than one community, then one should use InfoMap.
-If not, one should use standard modularity maximization with $\gamma=1$.
-That still leaves open the question of how the modularity should be maximized.
-In our studies we find the best results with simulated annealing, but simulated annealing is computationally expensive.
-In regimes where it is not feasible, we recommend using the Leiden algorithm instead.
-(Tests using other computationally efficient maximization schemes, such as the Louvain and spectral algorithms, generally performed less well than the Leiden algorithm.)
-IV Conclusions
-In this paper we have examined the performance of a range of mutual information measures for comparing labelings of objects in classification, clustering, or community detection applications.
-We argue that the commonly used normalized mutual information is biased in two ways: (1) because it ignores the information content of the contingency table, which can be large, and (2) because the symmetric normalization it employs introduces spurious dependence on the labeling.
-We argue in favor of a different measure, an asymmetrically normalized version of the so-called reduced mutual information, which rectifies both of these shortcomings.
-To demonstrate the effects of using different mutual information measures, we have presented results of an extensive set of numerical tests on popular network community detection algorithms, as evaluated by the various measures we consider.
-We find that conclusions about which algorithms are best depend substantially on which measure we use.
-Acknowledgements.
-The authors thank Samin Aref for useful comments and feedback.
-This work was supported in part by the US National Science Foundation under grants DMS–2005899 and DMS–2404617, and by computational resources provided by the Advanced Research Computing initiative at the University of Michigan.
-Appendix A Benchmarking
-In this appendix we give some additional details of our numerical tests.
-A.1 Results for the traditional symmetric normalized mutual information
-Figure 7: Performance of each of the six community detection algorithms considered here, as quantified by the conventional symmetrically normalized, non-reduced mutual information $I_{0}^{(S)}(c;g)$.
+As discussed in Section 2.5, all of these algorithms perform reasonably well, but the best performers in our tests are InfoMap and the two variants of modularity maximization.
+A.3 Results for the traditional symmetric normalized mutual information
+Figure 8: Performance of each of the six community detection algorithms considered here, as quantified by the conventional symmetrically normalized non-reduced mutual information $I_{0}^{(S)}(c;g)$.
 Figure 8 shows the performance of the same six community detection methods as in Fig.
-6, but measured using the standard, symmetrically normalized, non-reduced mutual information $I_{0}^{(S)}(c;g)$.
-By this measure many of the methods appear to perform implausibly well, far beyond the detectability threshold visible in Fig.
-6 in the regime where all methods should by rights fail.
-Note in particular the high scores achieved by the generalized modularity with $\gamma=10$ in this undetectable regime by virtue of the excessive number of groups it generates.
-A.2 LFR network generation
-The LFR networks we use for benchmarking are generated using the procedure described in [15], which is as follows.
+6, but measured using the standard, symmetrically normalized, non-reduced mutual information, which we denote by $I_{0}^{(S)}(c;g)$.
+By this measure, many of the methods appear to perform implausibly well, far beyond the detectability threshold visible in Fig.
+6, in the regime where all methods should by rights fail.
+Note in particular the high scores achieved by the generalized modularity with $\gamma=10$ by virtue of the excessive number of groups it generates.
+A.4 LFR network generation
+The LFR networks we use for benchmarking are generated using the procedure described in [20], which we summarize here.
   1. 1.
@@
 The parameter $\mu$ controls the relative number of edges within and between communities.
-For small $\mu$ there are many more edges within communities than between them, which makes the communities relatively easy to detect.
-But as $\mu$ gets larger there are more edges between communities and detection becomes more difficult in a manner reminiscent of the detectability threshold in the standard stochastic block model [30, 34].
+For small $\mu$ there are many more edges within communities than between them, which makes the communities easy to detect.
+But as $\mu$ gets larger there are more edges between communities and detection becomes more difficult in a manner reminiscent of the detectability threshold in the standard stochastic block model [21, 40], although there is no known analytic detectability threshold for the LFR model.
 In our studies we use values of $\mu$ in the range $[0.2,0.8]$.
@@
 Many networks are also found to have community sizes that approximately follow a power law, with typical exponents in the range from 1 to 2 [42, 43, 44, 20].
-We use $\tau_{2}=1.5$ with a minimum community size of $s_{\text{min}}=20$ in all cases, while the maximum community size is set to $s_{\text{max}}=\max\left\{n/10,100\right\}$.
+We use $\tau_{2}=1.5$ with a minimum community size of $s_{\text{min}}=20$ in all cases, while the maximum community size is set to $s_{\text{max}}=\max\left(n/10,100\right)$.
 Again, results were not particularly sensitive to these choices, provided they produce a valid distribution at all.
   4. 4.
-    Assign each node to a community randomly, one node at a time, while ensuring that the community chosen is always large enough to support the added node’s intra-community degree, given by $(1-\mu)k$ where $k$ is the total degree.
+    Assign each node to a community randomly, one node at a time, ensuring that the community chosen is large enough to support the added node’s intra-community degree, given by $(1-\mu)k$ where $k$ is the total degree.
   5. 5.
@@
 As in that study, we find that algorithm performance is dictated primarily by the parameters $n$ and $\mu$, so it is these parameters that are varied our summary figures.
-The LFR model is similar to a special case of the degree-corrected stochastic block model (DCSBM) [26], and hence one might expect that inference-based community detection methods employing the latter model would perform optimally on LFR networks.
+The LFR model is similar to a special case of the degree-corrected stochastic block model (DC-SBM) [36], and hence one might expect that inference-based community detection methods employing the latter model would perform well, perhaps even optimally, on LFR networks.
 Specifically, in the limit of an infinite number of sampled networks and perfect optimization of each community detection method, the final performance measure for any algorithm is given by the expectation value of the similarity $M(g,h[A])$ between the ground truth LFR partition $g$ and the partition $h[A]$ of the LFR network $A$ inferred using the algorithm, where the expectation is taken over the ensemble $P(A,g|\theta)$ of LFR networks and partitions $A,g$ generated using parameters $\theta$ (meaning $\mu$, $\tau_{1}$, $\tau_{2}$, etc.).
@@
 In other words, the optimal algorithm simply performs maximum a posteriori estimation under the model from which the network was generated.
-There is no explicit formula for the posterior probability under the LFR model, but to the extent that it is a special case of the DCSBM, we might expect the DCSBM (with appropriate priors) to give optimal results [39].
-The LFR model, however, is not precisely a special case of the DCSBM.
-In particular, the DCSBM normally assumes a uniform distribution over community sizes, where the LFR model assumes a power law.
-Moreover, we are not using the crude all-or-nothing error function: our entire purpose in this paper is to develop mutual information measures that aggregrate and weigh different modes of error in a sensible fashion.
-These differences, it appears, are enough to ensure that the DCSBM does not perform the best in our testing.
+There is no explicit formula for the posterior probability under the LFR model, but to the extent that it is a special case of the DC-SBM, we might expect the DC-SBM (with appropriate priors) to give optimal results [45].
+The LFR model, however, is not precisely a special case of the DC-SBM.
+In particular, the DC-SBM normally assumes a uniform distribution over community sizes, where the LFR model assumes a power law.
+Moreover, we are not using the crude all-or-nothing error function: our entire purpose in this paper is to develop mutual information measures that aggregate and weigh different modes of error in a sensible fashion.
+These differences, it appears, are enough to ensure that the DC-SBM does not perform the best in our testing.
 Regardless, we emphasize that our use of the LFR benchmark in our analysis is simply for consistency with previous studies of network community detection methods [20, 11, 13].
 The justification for our proposed similarity measure, on the other hand, is chiefly its theoretical merit over the conventional (symmetric, non-reduced) normalized mutual information, and is independent of the use of the LFR (or any other) benchmark.
+Appendix B Bias towards large numbers of groups in the traditional mutual information
+To shed light on why the traditional mutual information of Eq.
+(5) is biased towards an excessive number of groups and how the reduced mutual information of Eq.
+(4) corrects this, consider the (extreme) example in which every object is placed in a group on its own, producing a candidate labeling $c=(1,\ldots,n)$.
+Whatever the ground truth labeling $g$ is, this candidate labeling clearly provides no information about it whatsoever, so we expect the mutual information to be zero.
+But this is not what we find.
+The contingency table in this case is
+[…]
+so the conventional mutual information of Eq. (5) is
+$I_{0}(c;g)=\log\frac{n!}{\prod_{s}n_{s}^{(g)}!}=H_{0}(g).$ (16)
+This answer is not merely wrong; it is maximally so.
+The mutual information should take its minimum value of zero, but instead it takes the value $H_{0}(g)$, which is the maximum possible since $H_{0}(g)$ is an upper bound (see Eq.
+(2)).
+The reason for this result is that in this case the contingency table itself uniquely defines $g$, so neglecting it puts the mutual information in error by an amount equal to the complete information cost of the ground truth.
+If we include the the information cost of the table, on the other hand, as we do in the reduced mutual information, this erroneous behavior disappears.
+Assuming for simplicity that the ground-truth groups are of equal size, the optimal concentration parameters are $\alpha_{g|c}=0$ and $\alpha_{g}=\infty$, and the reduced mutual information of Eq.
+(4) becomes
+$I(c;g)$ $=I_{0}(c;g)-\log\left[\frac{n!}{\prod_{r}n_{r}!}(1/q_{g})^{n}\right]-n\log q_{g}$
+$=0.$ (17)
+which is the correct answer.
+Outside this special case, we observe from Eq. (4) that in the limit where the numbers of groups $q_{g}=q_{c}=q$ are held fixed and $n\rightarrow\infty$, the traditional mutual information grows as
+$I_{0}(c;g)\sim\mathrm{O}(n\log q)$ (18)
+while the correction term grows as
+$I(c;g)-I_{0}(c;g)\sim\mathrm{O}(q^{2}\log n).$ (19)
+Thus, in this limit the subleading term is relevant when $n/\log n\lesssim q^{2}/\log q$, a condition that holds in many practical contexts, where the number of communities can grow as $\sqrt{n}$ or faster.